### v1

# On the mechanics of inhaled bronchial transmission of pathogenic microdroplets generated from the upper respiratory tract, with implications for infection onset

Abstract
Could the microdroplets formed by viscoelastic stretching and break-up of mucosal liquids in the upper respiratory tract (URT), when inhaled further downwind through the airway, explain the brisk pace at which deep lung infections emerge following the onset of initial infection at the URT? While it is well-established that particulates inhaled from outside can possibly penetrate to the lower airway only if they are smaller than approximately 5 $\mathrm{\mu m}$, the fate of particulates (many exceeding 5-$\mathrm{\mu m}$ in diameter) that are sheared away from the intra-URT mucosa during inhalation remains an open question. These particulates predominantly originate in the region encompassing the nasopharynx, the oropharynx, and the laryngeal chamber with the vocal folds. To resolve the posed question, this study considers a computed tomography-based three-dimensional anatomical airway reconstruction and isolates the tract from the laryngeal vocal fold region (superior to the glottis), mapping the entire tracheal cavity and concluding at the second generation of the tracheobronchial tree in the lower airway. Through the delineated geometry, airflow simulation is conducted using the Large Eddy Simulation scheme to replicate relaxed inhalation at a rate of 15 L/min. Against the ambient air flux, representative numerical experiments have been performed to monitor the transport of liquid particulates (aerosols/droplets) with diameters ranging from $\mathrm{1~{}\mu m}$ to $\mathrm{30~{}\mu m}$, bearing physical properties akin to aerosolized mucus containing embedded virions. The full-scale numerically predicted transmission trends to the lower airway were found consistent with the findings from a reduced-order mathematical model that conceptualized the impact of intra-airway vortex instabilities on local particle transport through use of point vortex idealization in an anatomy-guided two-dimensional potential flow domain. The results collectively demonstrate markedly elevated trends of deep lung penetration by the URT-derived particulates, even if they are as large as 10 $\mathrm{\mu m}$ and 15 $\mathrm{\mu m}$. The high viral load physically carried by such droplets to the bronchial spaces could mechanistically explain the accelerated seeding of infection in the lungs.
preprint: APS/123-QED
I Introduction
When inhaled air sweeps past the mucociliary coating of the upper respiratory tract (URT), it is plausible to hypothesize that the interfacial interactions would lead to localized viscoelastic stretching and break-up of mucosal layers resulting in the formation and release of microdroplets, which would then be pushed downwind by the air streamlines. For instance, it is key to note the prominent instability effects when the viscoelastic layer (mucus) resting on a viscous fluid film (serous fluid) is exposed to the incoming airflow [1]. The ansatz is also supported by the established literature that has examined the reverse process: exhalation [2, 3, 4, 5]. The expiratory transport regimes, even during silent breathing, emit liquid particulates that span a wide range of length scales [6]. From such paradigm, if we now pivot our attention to inhaled (i.e., into the airway) transport, some immediate questions that come forth concern the fate of the intra-URT particulates generated during inhalation and their probable relevance in progressive disease transmission, especially to the deep lungs. On that note, the present study, through numerical experiments performed in a three-dimensional anatomical airway reconstruction and with simulation-informed reduced-order analytical modeling in an anatomy-inspired two-dimensional channel, attempts to answer the following main queries:
  1. $Q_{1}$.
    When inhaled from outside, it is well-known that the dominant inertial motion of the particulates larger than 5 $\mathrm{\mu m}$ ensures that hardly any of them, albeit potentially none at all, would penetrate to the lower airway. But is such conclusion also true for particulates of similar sizes generated through mucus separation from the URT surface during inhalation?
  2. $Q_{2}$.
    What is the viral load transmitted by particulates sheared away from the URT, that may manage to penetrate to the deep lungs, and more specifically, how does the viral transmission trend correlate with particulate sizes?
This study addresses $Q_{1}$ through full-scale numerical tracking of inhaled constituents inside an anatomically accurate airway geometry with a supporting reduced-order mathematical model of the system. Next for $Q_{2}$, it connects the fluid dynamics findings with virological parameters, such as the sputum viral concentration [7] for a representative pathogen (SARS-CoV-2), and critiques the projected viral load transmitted to the bronchial pathways in light of the verified infectious dose of the pathogen [8, 9]. The resulting translational analysis helps assess the proclivity for deep lung infection driven by inhalation of aerosolized intra-URT mucosal remnants. It is worth noting that this in silico approach, rooted in the underlying respiratory flow physics as discussed here, is agnostic to any virus specificity and is potentially extensible to other respiratory pathogens by integrating the fluid mechanics outcomes of inhaled transport with appropriate virological, epidemiological, and immunological data. In that context, this work could advance an emerging niche sub-field that brings together classical fluid dynamics and virology [10, 11, 12]. For select recent relevant studies on deep lung kinetics and pathophysiology, the reader may refer to [13, 14, 15, 16, 17].
Figure 1: (a) A demonstrative cartoon of the human respiratory system, encompassing the upper respiratory tract, the mouth, and the lower respiratory tract, extending till the deep lungs. The visual is adopted with copyrights’ permission from the Getty Images®. (b-c) Sample computed tomography (CT) imaging-based reconstruction of an adult human airway. It serves as a three-dimensional anatomically realistic equivalent of the cartoon in panel (a), with regions included till generation 2 (tertiary bronchi) of the tracheobronchial tree. The domain within the red box, i.e., panel (c), is isolated for the numerical experiments on inhaled downwind transport of microdroplets generated from the upper respiratory tract. The isolated region is additionally shown in panel (d) for the sagittal view with anatomical demarcations and in panel (e) for the coronal view. The symbol $g$ signifies the gravity direction in the numerical simulations and the subsequent analytical framework, with $x$, $y$, and $z$ defining the spatial orientation of the test cavity. Box (c) additionally highlights the precise length scale of the geometry.
II Methods
II.1 Numerical modeling of inhaled transport of microdroplets generated in the upper respiratory tract
As is the case for inhaled transport of microdroplets (also referred to as particulates, or equivalently, simply as particles in this exposition) through the respiratory cavity, in numerical simulations involving the dispersion of small particles under dilute conditions—the common approach assumes one-way coupling. This reduction implies that while the airflow continuum carries the particulates, the impact of such particulates on the underlying flow regimes could be disregarded. Consequently, the ambient airflow field is initially resolved, and the flow outcomes are subsequently employed to numerically solve the relevant particle equations of motion.
In this study, we have used Large Eddy Simulation (LES) scheme with dynamic subgrid-scale kinetic energy transport model to numerically replicate the inhaled air flux through an anatomically realistic airway reconstruction. To model the inhaled particle transport therein, we have applied the Lagrangian approach which is more suitable (compared to the Eulerian methods) for the dilute suspension of relatively large particles for which inertia could often be dominant in determining the spatial trajectories and eventual deposition spots. The simulated airflow field is coupled with the Lagrangian particle transport analysis to derive the intra-airway deposition and penetration trends.
II.1.1 Anatomical geometry reconstruction and spatial discretization
From de-identified, high-resolution, medical-grade computed tomography (CT) imaging, this study has first reconstructed a complete three-dimensional adult respiratory airway [18]; see Fig. 1a-c. The extraction warranted a radio-density thresholding between $-1024$ to $-300$ Hounsfield units [19, 20] to capture the airspace from the CT slices. Aiming to address $Q_{1}$ and $Q_{2}$ (posed in §I), the reconstruction has then been digitally redacted to focus on the space mapping the vocal fold region of the larynx (space with highest concentration of liquid particulates formed from intra-URT mucosal break-up) along with the distal laryngeal chamber, the trachea, and the lower airway extending till generation 2 of the tracheobronchial tree, with primary and tertiary bronchi, followed by the respiratory bronchiolar outlets as entry to the deeper recesses of the lungs. See Fig. 1d-e for the anatomical test geometry with the described regions marked out. To prepare the domain for numerical simulations, the test cavity was spatially segregated into 2.5 million graded, unstructured, tetrahedral elements, along with four layers of pentahedral cells (with 0.025-mm height for each cell and an aspect ratio of 1.1) extruded at the airway cavity walls [21] to resolve the near-wall particulate dynamics; see Fig. 2.
Figure 2: The black box in panel (a) marks the location of the demonstrated mesh region mapping a sagittal cross-section running mid-way through the test cavity (shown henceforth in Fig. 3). Panels (b) and (c) present multi-scale visuals of the mesh. In specific, panel (c) highlights the four layers of pentahedral cells used for near-wall refinement to ensure reliable numerical tracking of particle deposition.
While using the LES scheme, a finer mesh refinement typically leads to more precise outcomes. For instance, at the upper limit, achieving results akin to Direct Numerical Simulation (DNS) is possible if the grid dimensions are smaller than the Kolmogorov scale, $\mathcal{K}$, which is defined as follows [22]:
$\mathcal{K}=\left(\frac{\nu^{3}}{\varepsilon}\right)^{1/4},$ (1)
with $\varepsilon$ as the turbulence dissipation rate and $\nu$ being the fluid kinematic viscosity. Another length scale of note is the Taylor scale, $\lambda$, which typically exceeds $\mathcal{K}$ and is defined as
$\lambda=\left(\frac{10\nu k}{\varepsilon}\right)^{1/2},$ (2)
with $k$ being the turbulence kinetic energy. From the simulation data of the inhaled airflow field from this study, it is seen that both $\lambda$ and $\mathcal{K}$ collapse to $\mathcal{O}(10^{-4})$ m, while the mean grid scale is also $\rightarrow\mathcal{O}(10^{-4})$ m, suggesting that the test grid has been sufficiently resolved for precise estimation of the transport parameters.
II.1.2 Numerical simulation of inhaled airflow
We have employed the LES approach to numerically model the inhaled airflow, with eddies exceeding the grid scale explicitly resolved, whereas those that are smaller than the grid scale are approximated. Specifically, fluctuations below the grid size, referred to as subgrid scales, are filtered out, and their impact on larger scales is replicated through modeling. Assuming incompressible and isothermal conditions for the inhaled air flux, the filtered continuity and Navier–Stokes equations are respectively as follows:
$\frac{\partial}{\partial x_{i}}\left(\rho\overline{u}_{i}\right)=0$ (3)
and
$\frac{\partial\overline{u}_{i}}{\partial t}+\frac{\partial}{\partial x_{j}}\left(\overline{u}_{i}\overline{u}_{j}\right)=-\frac{1}{\rho}\frac{\partial\overline{p}}{\partial x_{i}}+\frac{\partial}{\partial x_{j}}\left(\nu\frac{\partial\overline{u}_{i}}{\partial x_{j}}\right)-\frac{\partial\tau_{ij}}{\partial x_{j}}.$ (4)
Here $\overline{u}_{i}$ represents the filtered (i.e., resolved) velocity, $\overline{p}$ is the filtered pressure, $\nu$ and $\rho$ are respectively the kinematic viscosity and the density of inhaled warmed-up air, and $\tau_{ij}$ is the subgrid scale (SGS) stress tensor defined by
$\tau_{ij}-\frac{1}{3}\tau_{kk}\delta_{ij}=-\nu_{sgs}\left(\frac{\partial\overline{u}_{i}}{\partial x_{j}}+\frac{\partial\overline{u}_{j}}{\partial x_{i}}\right),$ (5)
where $\nu_{sgs}$ is the SGS kinematic viscosity and $\delta_{ij}$ is the Kronecker delta. It is to be noted that $\tau_{kk}$, which comprises the isotropic part of the SGS stresses, is not modeled but added to the filtered static pressure. Subsequently, the instantaneous field velocity is given by
$u_{i}=\overline{u}_{i}+{u_{i}}^{sgs},$ (6)
with ${u_{i}}^{sgs}$ representing the SGS velocity fluctuations. The flow patterns and particle dispersion within the human respiratory system will be strongly impacted by the secondary flows common in such complex geometries and by the airflow transitions between laminar and turbulent regimes. To simultaneously capture the transitional features as well as the secondary flow formations, this study uses the dynamic subgrid-scale kinetic energy transport model [23, 24, 25]. Therein the SGS kinematic viscosity, $\nu_{sgs}$, is obtained from the Kolmogorov-Prandtl hypothesis [26] in the following form:
$\nu_{sgs}=C_{k}\,k^{1/2}_{sgs}\,\Delta_{f}.$ (7)
In the above equation, $C_{k}$ is a constant value and $\Delta_{f}$ is the filter size computed as $\Delta_{f}\equiv$ (grid cell volume)1/3. The term $k_{sgs}$ represents the SGS kinetic energy, defined by
$k_{sgs}=\frac{1}{2}\left(\overline{u_{i}u_{j}}-\overline{u}_{i}\overline{u}_{j}\right).$ (8)
To derive $k_{sgs}$, we solve the following filtered transport equation:
$\frac{\partial k_{sgs}}{\partial t}+\frac{\partial}{\partial x_{j}}\left(k_{sgs}\overline{u}_{j}\right)=\frac{\partial}{\partial x_{j}}\left(\nu_{sgs}\frac{\partial k_{sgs}}{\partial x_{j}}\right)+\frac{\partial\overline{u}_{i}}{\partial x_{j}}\left[\nu_{sgs}\left(\frac{\partial\overline{u}_{i}}{\partial x_{j}}+\frac{\partial\overline{u}_{j}}{\partial x_{i}}\right)-\frac{2}{3}k_{sgs}\delta_{ij}\right]-C_{\varepsilon}\frac{k^{3/2}_{sgs}}{\Delta_{f}},$ (9)
with the model constants in the previous equations, i.e., $C_{k}$ and $C_{\varepsilon}$, being determined dynamically [23].
Applying the LES scheme as described above, the inhaled airflow through the anatomical airspace was replicated for 15 L/min inhalation rate [27]. See Fig. 1d for the pressure inlet face at the location of vocal folds in the larynx and for the pressure outlets at the distal ends of the reconstructed generation 2 bronchial tubes. Enforcing no slip (i.e., zero velocity) boundary condition at the airway walls, the pressure gradient-driven simulation used time-steps of 0.0002 s [25], which is smaller than the Kolmogorov time scale = $\left(\nu/\varepsilon\right)^{1/2}$ [22], for a flow solution time of 0.35 s. The simulation was executed on a segregated solver with pressure-velocity coupling and second-order upwind spatial discretization. The eventual solution convergence was monitored by minimizing the mass continuity residual to $\mathcal{O}(10^{-3})$ and the velocity component residuals to $\mathcal{O}(10^{-6})$. Also, considering the warmed-up state of inhaled air passing through the respiratory pathway, the air density $\rho$ was set at 1.204 kg/m3 in the simulations, with 18.218$\times$10-6 m2/s as its kinematic viscosity $\nu$.
II.1.3 Numerical experiments for particulate transport
Against the solved airflow field, the particulates assumed to be formed from intra-URT mucus break-up were tracked from the vocal fold region of the larynx (see the marked inlet face in Fig. 1d). The air-particle phases were one-way coupled with the particles being impacted by the ambient flow field; the underlying flow domain was considered quasi-steady while evaluating the particle transport parameters. Lagrangian-based inert discrete phase model, with a Runge-Kutta solver, was used to numerically integrate the particle transport equation:
$\frac{du_{pi}}{dt}=\frac{18\mu}{d^{2}\rho_{p}}\frac{C_{D}Re_{p}}{24}\left(u_{i}-u_{pi}\right)+g_{i}\left(1-\frac{\rho}{\rho_{p}}\right)+F_{i}.$ (10)
Here $u_{pi}$ represents the particulate velocity, $\rho_{p}$ is the material density of the particulates, $d$ represents the particulate diameter, $Re_{p}$ is the particulate Reynolds number, $g_{i}$ signifies the gravitational acceleration in the $i$ direction, $C_{D}$ is the drag coefficient, and $F_{i}$ is any other additional body forces per unit particulate mass, e.g., the Saffman lift force exerted by a typical flow-shear field on small particulates transverse to the airflow direction. In this context, note that the solution scheme considered the particulates to be large enough to ignore any Brownian motion effects on their dynamics.
To evaluate the drag term (i.e., the co-factor on the first term) on the right hand side of equation 10, the quantities $Re_{p}$ and $C_{D}$ are respectively computed as
$Re_{p}$ $=\frac{\rho_{p}d\left|u_{i}-u_{pi}\right|}{\mu}$ $C_{D}$ $=a_{1}+\frac{a_{2}}{Re_{p}}+\frac{a_{3}}{{Re_{p}}^{2}},$ (11)
where $\mu$ is the molecular viscosity of the ambient fluid (i.e., air), while $a_{1}$, $a_{2}$, and $a_{3}$ are functions of $Re_{p}$ determined based on the spherical drag law [28]. Subsequently, the particulate trajectories are derived from their spatiotemporal locations, $x_{i}\left(t\right)$, obtained by numerical integration of the following velocity vector equation:
$u_{pi}=\frac{dx_{i}}{dt}.$ (12)
Intra-airway aerial tracking of a particulate pathline was stopped once it entered the mesh element layer adjacent to the enclosing walls of the respiratory cavity. The particulates escaping through the bronchiolar outlets were also recorded, and represented the particulates penetrating to the respiratory bronchioles and deep lungs. The numerical experiments in this study tested particulates of diameters $\mathrm{1-30~{}\mu m}$ (with increments of 1 $\mathrm{\mu m}$). In total, $\mathcal{N}$ = 1622 particulates of each size were tracked; $\mathcal{N}$ being the number of surface elements in the mesh layer mapping the inlet face (see Fig. 1). Also, considering that saliva-mixed mucus is 99.5% water [29], the material density of the particulates was assumed to be $\rho_{p}$ = 1.00175 g/ml, approximated as the weighted average (see equation 13 below) between 99.5% water and the residual pathogenic and biological non-volatile compounds with a representative density of $\rho_{nv}=$ 1.35 g/ml; e.g., for protein [30], the density being independent of the nature of the protein and particularly independent of its molecular weight. The weighted average calculation for the simulated particulate density is as follows:
$\rho_{p}=\frac{0.5\times\rho_{nv}+99.5\times\rho_{w}}{100},$ (13)
with $\rho_{w}=1.0$ g/ml representing the density of water. The consideration of inertness for the tracked particulates (per equation 10) also implies that the modeling framework was agnostic to the biological nuances of the embedded constituents in the particulates, beyond the imposition of the appropriate physical properties, e.g., density. The modeling approach also discounted any heat transfer effects between the flow constituents and the surrounding tissues enclosing the anatomical airspace.
Figure 3: (a) Laryngotracheal region upwind from the primary bronchi used to develop a simulation-informed reduced-order analytical model (S-ROAM) for bronchial transmission. AA marks the distal larynx upwind face and BB marks the downwind face of the distal tracheal cavity; see labels in Fig. 1d. The cross-sectional areas of the cavity at AA and BB are 108.61 mm2 and 123.03 mm2, respectively. The corresponding hydraulic diameters are 11.03 mm and 12.20 mm, respectively. The linear spatial distance between the two faces is approximately 115.84 mm, at angle $\theta_{1}$ = 23.77 to the vertical (direction of gravity in the numerical simulations). Panel (b) highlights the location and orientation of the planar cross-section shown in panel (c) and also previously as the location of the mesh visuals in Fig. 2. The vortex strengths and positions for the S-ROAM are extracted from the simulated data mapping this two-dimensional plane; the plane cuts through the entire cavity of panel (a). $\theta_{2}$ = 33.94 is the assumed angle in S-ROAM, at which microdroplets enter the AA inlet face and is governed by the anatomical shape of the cavity upwind from AA (marked by the dashed black traces; also see Fig. 1).
II.2 Analytical model for inhaled airflow and particulate transport
To verify the particulate transport trends to the bronchial domain as derived numerically, this study invokes a simulation-informed reduced-order analytical model (S-ROAM), with a two-dimensional channel mimicking the laryngotracheal domain. Therein, the inhaled airflow was modeled in the dimensionless complex $\chi$ plane, with $\chi=\alpha+\mathrm{i}\,\beta$ and $\mathrm{i}\,\beta$ aligned with the channel’s streamwise axis; see Fig. 3. The two-dimensional channel had its streamwise length and cross-stream width based on the averaged dimensions of the anatomical tract from Fig. 3a-b. The channel walls were placed at $\alpha=0,1$. Based on the geometric inputs, the S-ROAM enforced $L/W\approx 9$, where $L$ represents the streamwise channel length and $W$ is the inlet width. For additional details on the analytical approach, see our recent paper [31].
The reduced system modeled an extended vortex patch emerging in the full-scale numerical simulation as a collection of five point vortices distributed over the area of the simulated vortex core (on the channel plane, mapping the mid-sagittal section of the three-dimensional cavity; see Fig. 3b), bearing dimensionless circulations $\Gamma_{i}=\omega A/5UW$, where $\omega$ was the mean vorticity (as determined from the numerical simulations) spread over the patch area $A$ and $U$ is the characteristic streamwise air speed through the channel. From the simulated data, we had: $\omega=1750$ s-1, $A=2.65\times 10^{-5}$ m2, and $U=Q/a=2.16$ m/s, with $a=$ 115.82 mm2 being the area of the inlet face and $Q=15$ L/min $=0.00025$ m3/s being the simulated inhalation flux. The anatomical planar section from which the vorticity information was extracted is shown in Fig. 3c.
In the S-ROAM, the straight streamlines at the channel walls were established through inclusion of periodic images of the point vortices in the cross-stream $\alpha$ direction. The resulting (dimensionless) complex potential for this system, with a dimensionless background speed of unity, is [32, 33, 34]:
$F(\chi)=\phi(\alpha,\beta)+\mathrm{i}\,\psi(\alpha,\beta)=\mathrm{i}\,\chi+\sum_{i=1}^{5}\frac{\Gamma_{i}}{2\pi\mathrm{i}}\log\left\{\frac{\sin\left[\pi(\chi-\chi_{i})/2\right]}{\sin\left[\pi(\chi+\chi_{i}^{*})/2\right]}\right\},$ (14)
with $\phi$ as the velocity potential and $\psi$ as the real-valued flow streamfunction; the asterisk denotes complex conjugation.
Subsequently, the inhaled particulate motion was analytically derived using a simplified version of the Maxey-Riley equation [35], in the two-dimensional vector form [36, 37]:
$\frac{d{\mbox{$w$}}}{dt}=-\left[{\mbox{$J$}}+\frac{2\,{St}^{-1}}{3(\sigma+1)}{\mbox{$I$}}\right]\cdot{\mbox{$w$}}+\frac{\sigma}{\sigma+1}\left({Fr}^{-2}{\mbox{$g$}}-\frac{\text{D}{\mbox{$u$}}}{\text{D}t}\right).$ (15)
[…]
Although viscous effects are neglected in the above velocity model, particle drag is included. This form of the model ignores the Faxen correction terms and the Basset-Boussinesq history force. Particulates are assumed to be entering the S-ROAM domain (inclined at an angle $\theta_{1}=23.77^{\circ}$ to the direction of gravity, per the general shape and orientation of the test anatomical cavity) with speed $U$ and at an angle of $\theta_{2}=33.94^{\circ}$ with respect to the S-ROAM’s $\alpha$-axis, mimicking the simulated motion of particles entering the region; see Fig. 3c. As representative examples, particulate trajectories were derived for two initial positions at the mid-point and near the right edge of the reduced channel’s inlet span, and for seven different particulate diameters, namely, 5, 10, 15, 20, 25, 30, and (as an extremal case) 50 $\mathrm{\mu m}$.
II.3 Connecting the fluid dynamics outcomes to virology
From the numerical experiments, this study has deduced the simulated deposition and penetration efficiencies of the particulates along the bronchial tubes (see Table 1) and evaluated the respective volumetric transmission to the bronchi. The projected net deposited (at the primary and tertiary bronchi) and penetrated (moving into the respiratory bronchioles) volumes are then multiplied with the sputum viral concentration for a specific pathogen (in this study, SARS-CoV-2) to evaluate and compare the viral load transported via inhaled aerial advection of URT-derived liquid particulates to the lower airway and deep lungs, for select test particulate sizes, namely, 1, 5, 10, and 15 $\mathrm{\mu m}$.
Figure 4: 30 representative velocity streamlines extracted from the numerically simulated inhaled airflow field.
Note that the average sputum viral concentration for SARS-CoV-2 has been reported as $\mathcal{V}$ = 7.0$\times$106 virions/ml, through count measurements of the RNA copies for the single-stranded virus present in the airway liquid samples collected from hospitalized COVID-19 patients [7].
III Results
Table 1: Numerically simulated bronchial deposition and penetration data: The particulate sizes, that have been explored further in Table 2 (to formalize the viral transmission trends as a function of the microdroplet dimensions), are in bold font. Symbols: $d$ = tested aerosol (or, droplet) diameter; $\mathcal{N}$ = total number of aerosols/droplets tracked for each particulate diameter; $n_{p}$ = number of deposited particulates in the primary bronchi; $n_{t}$ = number of deposited particulates in the tertiary bronchi; $n_{d}$ = number of particulates penetrating into the respiratory bronchioles toward the deep lungs; $\eta$ = deposition (or, penetration) rate (in %) to the bronchial pathways.
Separated aerosol/droplet size
($d$, in $\mathrm{\mu m}$)
$\mathcal{N}$
$n_{p}$
$n_{t}$
$n_{d}$
Lung penetration ($\eta$, in %)
1 1622 11 0 1611 100.00
2 1622 4 0 1617 99.94
3 1622 14 0 1608 100.00
4 1622 16 0 1606 100.00
5 1622 32 3 1587 100.00
6 1622 83 13 1526 100.00
7 1622 123 25 1469 99.69
8 1622 170 41 1381 98.15
9 1622 173 34 1291 92.36
10 1622 210 41 1183 88.41
11 1622 269 47 1054 84.46
12 1622 316 43 933 79.65
13 1622 369 47 797 74.78
14 1622 420 88 607 68.74
15 1622 461 93 433 60.85
16 1622 489 89 293 53.70
17 1622 460 91 217 47.35
18 1622 417 99 144 40.69
19 1622 365 91 80 33.05
20 1622 324 78 37 27.07
21 1622 280 59 17 21.95
22 1622 221 48 12 17.32
23 1622 172 17 10 12.27
24 1622 120 4 1 7.71
25 1622 81 3 0 5.18
26 1622 57 1 0 3.58
27 1622 38 0 0 2.34
28 1622 20 0 0 1.23
29 1622 5 0 0 0.31
30 1622 0 0 0 0.00
III.1 Bronchial deposition and deep lung penetration as projected from numerical modeling
Figure 4 presents sample inhaled airflow velocity streamlines from the numerical simulation of inhaled airflow passing through the CT-reconstructed anatomical cavity. The simulated inhalation rate of 15 L/min to mimic relaxed breathing warranted an inlet-to-outlet pressure gradient of -9.5 Pa. Against the flow field, the test particulates were tracked with their starting locations uniformly distributed on the cross-sectional space spanning the vocal fold region of the laryngeal cavity (commensurate with their hypothesized formation sites at the URT—through break-up of mucus strata along the nasopharynx, the oropharynx, and the vocal folds).
Table 1 details the lower airway deposition and penetration data for the tested particulates bearing diameters $\mathrm{1-30~{}\mu m}$ (with increments of $\mathrm{1~{}\mu m}$). While it is expected that the particulates $\lesssim$ $\mathrm{5~{}\mu m}$ would relatively comfortably penetrate to the deep lungs (as is clearly the case per Table 1), the high transmission percentages, e.g., for even the 10- and 15-$\mathrm{\mu m}$ particulates is striking—them being 88.41% and 60.85%, respectively. In the table (for 1622 particulates tracked for each size), $n_{p}$ quantifies the number of particulates depositing in the primary bronchi, $n_{t}$ quantifies the number of particulates depositing at the tertiary bronchi, and $n_{d}$ measures the number of particulates escaping through the respiratory bronchiolar outlets of the in silico model and subsequently moving into the deeper recesses of the lungs. See Fig. 1 to review the anatomical zone nomenclature for this study and the respective locations of the different physiological regions along the airway.
III.2 Viral load transmitted to the lower airway
Table 2 lists the viral load transmitted to the bronchial pathways, for the representative test particulate sizes 1, 5, 10, and 15 $\mathrm{\mu m}$. For the time scale $T$ (in hours) over which the viral load transmission is to be estimated, we have used 3 days (i.e., 72 hours) based on reported data [38] on the typical time interval that has revealed confirmed infection onset in the deep lungs subsequent to the emergence of initial symptoms along the URT. Further, applying $\delta t$ = 5 s as the average duration for a complete breathing cycle [39], the viral load, $L_{v}$, transmitted to the bronchial pathways and deep lungs can be computed as:
$L_{v}=\frac{144\,\pi\,d^{3}\,\eta\,T\,\mathcal{V}\,\mathcal{P}\left(d_{i}\right)}{\delta t}\times 10^{-12}.$ (16)
where $\mathcal{P}\left(d_{i}\right)$ is the particle size density function that quantifies the number of microdroplets of size $d_{i}$ generated during each inhalation cycle. For simplicity, the data in Table 2 enforced $\mathcal{P}$ = 1, irrespective of the particulate size, i.e., the particulate formation rate was conservatively assumed to be 1 per breathing cycle, for each test particulate size. Finally, note that while the reported numbers in Table 2 are based on the estimates guided by equation 16, the zero viral load assessment for the 1-$\mathrm{\mu m}$ particulates implies that the corresponding transmission estimate from equation 16 resulted in a fractional number $\ll$ 1, and bears no physical significance.
Table 2: Magnitude of viral load transmission into bronchial pathways as a function of inhaled particulate sizes: The duration in consideration is 3 days, with a conservative estimate of 1 particulate of each test size generated during each breathing cycle. The breathing cycles last 5 s [39].
Separated aerosol/droplet size
($d$, in $\mathrm{\mu m}$)
Number of virions ferried to the bronchial pathways
($L_{v}$)
15 390
10 168
5 24
1 0
Figure 5: Log-log plot for $\eta$, summing up the bronchial deposition and deep lung penetration percentages (see Table 1), for each numerically tracked microdroplet size, $d$. The simulation-derived data points are marked by the dark blue circles. The solid black line is a fitted curve of the Heaviside step function form (see equation 17).
III.3 Heaviside trend governing bronchial deposition and penetration
Collapsing the net particulate transmission efficiencies from Table 1, labeled as $\eta$ (in %), on to a log-log scale with particulate diameters $d$ (in $\mathrm{\mu m}$) placed along the horizontal axis, reveals (as is expected) a high lower airway penetration for smaller particulate sizes, followed by a gently sloping dip; see Fig. 5. The behavior could be approximately mathematized with a modified Heaviside function [40] of the following analytical form:
$\eta(d)=\mathcal{C}_{1}+e^{\mathcal{C}_{2}/\left[1+e^{2k\left(d+\mathcal{C}_{3}\right)}\right]},$ (17)
where $k$ and $\mathcal{C}_{i}$, with $i\in\{1,2,3\}$, are constant fitting parameters. For the solid black curve plotted in Fig. 5 and fitting through the data-points from the numerical experiments, the following values were used:
$\mathcal{C}_{1}=-1.95,\quad\mathcal{C}_{2}=4.95,\quad\mathcal{C}_{3}=-25.0,\quad k=0.08.$ (18)
Figure 6: (a) Simulated vorticity contour map on a representative two-dimensional cross-section (the S-ROAM domain) running approximately midway through the cavity; see Fig. 3b. The region with the dominant vortex core is marked by $\mathbb{V}$. (b-h) show two sample particle pathlines (in red) against the S-ROAM streamlines (in grey), respectively for particle sizes 5, 10, 15, 20, 25, 30, and 50 $\mathrm{\mu m}$. The black lines on either side of the S-ROAM domain mark the slip wall boundaries. In each case, the vortex patch $\mathbb{V}$ is mimicked by a set of five point vortices embedded on a the two-dimensional flow field with background unidirectional flow in the $\beta$-direction. The representative particle entering near the middle of the inlet face is marked as $\mathbb{I}$ and the particle entering near the right edge of the model channel is marked as $\mathbb{II}$.
III.4 Analytical projections for particulate transport: consistent with the full-scale numerical findings
Figure 6a demonstrates the vorticity field mapped over the section shown in Fig. 3c. In the reduced-order analytical setup, the vortex patch marked by $\mathbb{V}$ is modeled with a spatial assembly of five point vortices, as described in §II.2. In panels b-h of Fig. 6, the grey curves show the streamlines of the background flow field in the S-ROAM, with the red curves tracing the sample particulate trajectories. Smaller particulates, owing to their low inertia and smaller $St$, follow the streamlines (on which they were embedded at entry points) more closely. For larger particulates, the inertia-dominated dynamics would inhibit downwind penetration through biasing their motion toward wall deposition, e.g., see panels g and (especially) h.
Note that the particulates entering through the middle of the inlet face of the S-ROAM are marked as $\mathbb{I}$ while those entering near the right edge are marked as $\mathbb{II}$. As shown in Fig. 7, the particulates $\mathbb{II}$, being nearer to the vortex region, are deviated more (from the airflow streamlines they were embedded on at the channel inlet) compared to the particulates $\mathbb{I}$. The deviation generically grows as the particulate diameters are increased, which aligns with the findings in Table 1. A greater deviation implies that the particulates are being increasingly shifted toward the channel walls, which in turn would result in a declining deposition and penetration levels in the lower airway.
Figure 7: Comparison of the absolute deviation ($\tilde{\mathbb{D}}$, normalized with respect to the S-ROAM channel width) of the particle pathlines from the respective streamline they were embedded on at the inlet face. The normalized deviation is measured along the $\alpha$-direction (see Fig. 6). Deviation curves for particles entering near the middle of the S-ROAM inlet face is marked as $\mathbb{I}$ and for the particles entering near the right edge of the model channel is marked as $\mathbb{II}$.
IV Open questions: Perspectives on enhancing the biophysical realism of the modeling approach
IV.1 On the interfacial mechanics at mucociliary layers during inhalation
While the numerical and analytical models presented here conclude that aerosols/droplets of the length scale $\mathcal{O}$($10^{1}$) $\mathrm{\mu m}$, formed through shearing and disintegration of intra-URT mucosal filaments, can have significant deposition and penetration in the deep lungs (e.g., see Table 1); to enhance the physiological realism of the model—it is essential that we obtain the actual size distribution and formation rate of liquid particulates generated through viscoelastic separation as the inhaled air brushes past the URT mucus. The present study employed the conservative estimate that 1 particulate of each tested size is formed during each breathing cycle and is a limitation. The reader should however note that unlike expulsion regimes (during exhalation), where imaging-based data collection is relatively straightforward with human subjects expelling particulates into the outside air for different speech and breathing parameters, the current problem of characterizing the internal reverse transport (inhalation) into the lower airway could be somewhat challenging.
Figure 8: (a) Nasopharynx (compare with labels on Fig. 1), which is the initial trigger site for SARS infection owing to the propensity of ciliated epithelial cells with angiotensin-converting enzyme 2 (ACE2) surface receptors that the spike protein viruses bind to during cell intrusion. (b) Representative planar outline of the surface topology at the nasopharynx. Phase 1 stands in for the inhaled air flux, phase 2 signifies the mucus film. The blue arrow is a cartooned depiction of the inhaled airflow profile sweeping over the mucosal topography (in pale red). The black curve in panel (b) delineating the mucosal surface topography is extracted from CT data.
With in vivo tests in human subjects all but unlikely to receive institutional review board approvals, the approach could be essentially two-fold with synergistic numerical modeling and experimental visualizations. One can consider a 2-phase interaction on anatomically realistic upper airway surfaces; the two phases being the mimicked versions of inhaled air (phase 1) and the relatively static mucosal film (phase 2); see Fig. 8. The interfacial mechanics leading to particulate formation and release, along with their spatiotemporal growth and size distribution, could then be assessed with multiphase simulation techniques. While similar problems in a different context have been studied before—the three-dimensional wavy anatomic topology of the URT surface can potentially reveal interesting nuances related to the dynamics. Experimental visualizations with air flux brushing past a viscoelastic film mimicking mucus rheology should then be able to validate the micro-scale mechanics revealed by the computational approach.
IV.2 On experimental validation of deep lung penetration through in vitro physical tests
To verify the numerical modeling of the overall spatial transport, experiments could be conducted in 3D-printed anatomically realistic airway casts that are internally coated with (say) concentrated glycerol (with tuned dilution levels) standing in for mucus. Controlled air flux, with embedded aerosols, could be passed through the cast via suction mechanism set up with a vacuum pump. One approach (amongst others) for getting highly-resolved data on the lower airway penetration could be to use the sophisticated gamma scintigraphy technique [41], wherein the aerosolizing solution to be administered into the three-dimensional cast (with the incoming air) would be seeded with a mildly radioactive element (e.g., Technetium). After the particulates have landed along the airway walls, the radioactive signals (emitted by the deposited mass and hence presenting a measure of local penetration) could be compared with the in silico deposition patterns along the bronchial pathways. Note that for this approach to be functionally consistent, the printed cast should be separable into (at least) two parts along the sagittal mid-section to facilitate controlled liquid coating of its internal surfaces. The separable 3D-printed pieces will also logistically facilitate radio-signal imaging of the wall deposits.
V Conclusion: the main inferences
V.1 Can large particulates, generated from the intra-URT mucus coating during inhalation, penetrate to the bronchial airspace?
The particulate transmission trends (see Table 1) from the numerical experiments performed in this study are found to be consistent with the first-principles reduced-order analytical findings that modeled the impact of dominant intra-airway vortex instabilities in the laryngotracheal domain on local particle transport (Fig. 6). Clearly, the results establish that not only the aerosols (i.e., particulates with diameters $\lesssim$ 5 $\mathrm{\mu m}$) but also droplets as large as 10 and 15 $\mathrm{\mu m}$ can exhibit remarkable efficiency at reaching the bronchial spaces and deep lungs, so long as they are sheared away from the URT surface and are still airborne as they enter the larynx. This stands in contrast to the general perception that only the particulates that are smaller than 5 $\mathrm{\mu m}$ can penetrate to the lower airway. Missing the nuance therein is the fact that the incumbent perspective is based on the mechanics of particulates inhaled from the outside air, and our understanding of intra-URT mucosal break-up during inhalation and the subsequent aerial advection profiles in the downwind tract is still nascent.
Figure 9: (a) Sagittal cutting plane through the anatomical airway to extract the intersected curve shown in panel (b) The latter is used to evaluate the tortuosities between P-R and Q-R tracts of the airway.
For a physics-based rationale to explain the derived deposition and penetration profiles, let us consider the tortuosity $\mathcal{T}$ of the pathways to be traversed by a particulate inhaled from outside (called hereafter $P_{o}$), compared to a particulate that is generated from the URT and is still airborne in the laryngeal airspace (hereafter called $P_{u}$). Figure 9 illustrates a representative vertical cutting plane used to extract the tortuosity measurements. For $P_{o}$, the corresponding tortuosity is the ratio of the curved path length and the linear distance in space between the points P and R; let us represent it as $\mathcal{T}_{P_{o}}$. For $P_{u}$, it is similarly the ratio of the curved path length and the linear distance in space between the points Q and R; let us represent it as $\mathcal{T}_{P_{u}}$. The geometric measurements of the spatial distances return: $\mathcal{T}_{P_{o}}\approx 1.90$ and $\mathcal{T}_{P_{u}}\approx 1.04$. Consequently, if the mechanics of $P_{o}$ is inertia-dominated (true for particulates $\gtrsim$ 5 $\mathrm{\mu m}$), they exhibit less success at navigating the highly tortuous pathway (which would have otherwise led them to the lower airway) and are deposited along the anterior URT. On the contrary, particulates of similar sizes, if they only have to traverse the Q-R tract (see Fig. 7), the less tortuous pathway ensures that a high percentage of them would end up depositing along (or, penetrating to) the bronchial domains. As an aside, also note that the tortuosity estimates obtained here match exactly with our previously published data on mammalian airway morphology with different test airway cavities [42]; see panel (d) in Fig. 1 of the cited article. Thus, the findings from this study satisfactorily address the question $Q_{1}$ posed in the introduction (§I), a variant of which is the question heading this subsection.
V.2 How do the transmitted viral loads compare to the infectious dose of the test pathogen?
Infectious dose, $I_{D}$, of a virus quantifies the minimum number of virions that can potentially launch infection in an exposed subject [43, 29] and is a fundamental virological parameter. Independent studies by us [8, 44] and others [9, 45] have verified that $I_{D}\approx$ 300, for SARS-CoV-2 onset. Thus, evidently (per Table 2), the viral load transmitted by, e.g., the 15-$\mathrm{\mu m}$ droplets would alone exceed the $I_{D}$ threshold, thereby providing a mechanics-based rationale for the fast disease progression to the lower airway. The brisk pace is otherwise difficult to explain based exclusively on tissue level proliferation and direct deep lung inhalation of sub 5-$\mathrm{\mu m}$ particulates from outside. However, from a translational paradigm, it is critical to note that the present in silico framework does not yet take into account the host innate and adaptive immune responses to the invading virions. The immunological considerations, once incorporated into the mechanics paradigm, can help rationalize the varying rates of clinical prognosis recorded in different subjects [38]; e.g., the deep lung infection for SARS-CoV-2 has historically ensued over a range of $\mathrm{2-8}$ days in different subjects following the appearance of initial symptoms along the upper respiratory tract. In summary, the findings do satisfactorily address the question $Q_{2}$ posed in §I, a variant of which is the question heading this subsection.
Acknowledgments
The author thanks the National Science Foundation for support via the NSF CAREER Grant, Award No. CBET 2339001 (Fluid Dynamics program). The author also thanks Neelesh Patankar (Northwestern University) for stimulating discussions on bronchial fluid mechanics, Julia Kimbell (School of Medicine at the University of North Carolina Chapel Hill) for providing access to existing, de-identified CT imaging, and Abby Wortman (Clinical and Lab Instructor for the Respiratory Care program at the author’s institution) for checking the nomenclature on airway physiology used in this study.
Data sharing
Supplemental information (including simulated data, raw files, and codes) are available on-request via OneDrive.

### v2

@@
 Abstract
-Could the microdroplets formed by viscoelastic stretching and break-up of mucosal liquids in the upper respiratory tract (URT), when inhaled further downwind through the airway, explain the brisk pace at which deep lung infections emerge following the onset of initial infection at the URT?
-While it is well-established that particulates inhaled from outside can possibly penetrate to the lower airway only if they are smaller than approximately 5 $\mathrm{\mu m}$, the fate of particulates (many exceeding 5-$\mathrm{\mu m}$ in diameter) that are sheared away from the intra-URT mucosa during inhalation remains an open question.
-These particulates predominantly originate in the region encompassing the nasopharynx, the oropharynx, and the laryngeal chamber with the vocal folds.
-To resolve the posed question, this study considers a computed tomography-based three-dimensional anatomical airway reconstruction and isolates the tract from the laryngeal vocal fold region (superior to the glottis), mapping the entire tracheal cavity and concluding at the second generation of the tracheobronchial tree in the lower airway.
-Through the delineated geometry, airflow simulation is conducted using the Large Eddy Simulation scheme to replicate relaxed inhalation at a rate of 15 L/min.
-Against the ambient air flux, representative numerical experiments have been performed to monitor the transport of liquid particulates (aerosols/droplets) with diameters ranging from $\mathrm{1~{}\mu m}$ to $\mathrm{30~{}\mu m}$, bearing physical properties akin to aerosolized mucus containing embedded virions.
-The full-scale numerically predicted transmission trends to the lower airway were found consistent with the findings from a reduced-order mathematical model that conceptualized the impact of intra-airway vortex instabilities on local particle transport through use of point vortex idealization in an anatomy-guided two-dimensional potential flow domain.
-The results collectively demonstrate markedly elevated trends of deep lung penetration by the URT-derived particulates, even if they are as large as 10 $\mathrm{\mu m}$ and 15 $\mathrm{\mu m}$.
-The high viral load physically carried by such droplets to the bronchial spaces could mechanistically explain the accelerated seeding of infection in the lungs.
+Could the microdroplets formed by viscoelastic fragmentation of mucosal liquids within the upper respiratory tract (URT) explain the brisk onset of deep lung infection following initial URT infections?
+Generally, particulates, inhaled through the nostrils and therefore navigating the intricate topography of the anterior nasal cavity, can efficiently reach the lower airway only if they are small enough, typically $\lesssim 5\,\mathrm{\mu m}$.
+However, the fate of larger particulates, many exceeding 5-$\mathrm{\mu m}$ in diameter, that are sheared from the initial infection sites along the intra-URT mucosa during inhalation remains unresolved.
+These particulates originate primarily from the nasopharynx, oropharynx, and the laryngeal chamber containing the vocal folds.
+To investigate, this study employs a computed tomography-based three-dimensional anatomical airway reconstruction, isolating the tract from the larynx and mapping the tracheal cavity through to the third generation of the tracheobronchial tree; constituent transport across the distal bronchial outlets is also recorded to assess deep lung penetration.
+Within the defined geometry, airflow simulations are conducted with the Large Eddy Simulation scheme to replicate relaxed inhalation at 15 L/min flow rate.
+Against the ambient air flux, numerical experiments are performed to monitor the transport of particulates (aerosols/droplets) with diameters $\mathrm{1-30~{}\mu m}$, bearing physical properties akin to aerosolized mucus with embedded virions.
+The full-scale numerical transmission trends are consistent with findings from our reduced-order mathematical model that conceptualizes the influence of intra-airway vortex instabilities on local particle transport through point vortex idealization in an anatomy-guided two-dimensional potential flow domain.
+The results collectively demonstrate markedly elevated lower airway penetration by URT-derived particulates, even by those as large as 10 and 15 $\mathrm{\mu m}$.
+The high viral load, often exceeding the pathogen-specific infectious dose, carried by such droplets into the bronchial spaces of the sample airway, provides a plausible mechanistic explanation for the accelerated seeding of secondary lung infection.
 preprint: APS/123-QED
 I Introduction
-When inhaled air sweeps past the mucociliary coating of the upper respiratory tract (URT), it is plausible to hypothesize that the interfacial interactions would lead to localized viscoelastic stretching and break-up of mucosal layers resulting in the formation and release of microdroplets, which would then be pushed downwind by the air streamlines.
-For instance, it is key to note the prominent instability effects when the viscoelastic layer (mucus) resting on a viscous fluid film (serous fluid) is exposed to the incoming airflow [1].
-The ansatz is also supported by the established literature that has examined the reverse process: exhalation [2, 3, 4, 5].
-The expiratory transport regimes, even during silent breathing, emit liquid particulates that span a wide range of length scales [6].
-From such paradigm, if we now pivot our attention to inhaled (i.e., into the airway) transport, some immediate questions that come forth concern the fate of the intra-URT particulates generated during inhalation and their probable relevance in progressive disease transmission, especially to the deep lungs.
-On that note, the present study, through numerical experiments performed in a three-dimensional anatomical airway reconstruction and with simulation-informed reduced-order analytical modeling in an anatomy-inspired two-dimensional channel, attempts to answer the following main queries:
+When inhaled air sweeps past the mucociliary coating of the upper respiratory tract (URT), interfacial interactions lead to localized viscoelastic stretching and breakup of mucosal layers resulting in the formation and release of microdroplets [1, 2, 3], which could then be pushed downwind by the airflow streamlines.
+Prominent instability effects when the viscoelastic layer (mucus) resting on a viscous fluid film (serous fluid) is exposed to the incoming airflow are to be noted in this context; e.g., see [4].
+The mechanism is analogous to numerous observations from the reverse process: exhalation [5, 6, 7, 8].
+The expiratory transport regimes, for intense respiratory events [9] and even during silent breathing [10], often involve mucus fragmentation and subsequent emission of liquid particulates spanning a wide range of length scales.
+From that paradigm, if we pivot our attention to inhaled (i.e., into the airway) transport, some immediate questions that come forth concern the fate of the intra-URT particulates generated during inhalation and their probable relevance in progressive disease transmission, especially to the deep lungs.
+On that note, the present study, through numerical experiments performed in a three-dimensional anatomical airway reconstruction and with simulation-informed reduced-order analytical validation in an anatomy-inspired two-dimensional channel, attempts to answer the following main queries:
   1. $Q_{1}$.
-    When inhaled from outside, it is well-known that the dominant inertial motion of the particulates larger than 5 $\mathrm{\mu m}$ ensures that hardly any of them, albeit potentially none at all, would penetrate to the lower airway.
-But is such conclusion also true for particulates of similar sizes generated through mucus separation from the URT surface during inhalation?
+    If inhaled from outside and consequently navigating the complex anterior nasal topography, the dominant inertial motion as well as gravitational impaction for particulates larger than 5 $\mathrm{\mu m}$ may, in general, prevent their penetration to the lower airway [11, 12].
+However, is this also true for particulates of similar sizes shed through mucus separation at the back of the nasal passage from the URT sub-sites like the nasopharynx, oropharynx, and from around the vocal folds during inhalation?
+Hypothetically, with such particulates still being airborne at the larynx (see Fig 1), the relatively straight structural shape of the downwind tract up to the tracheal base may facilitate their penetration into the lower airway.
   2. $Q_{2}$.
-    What is the viral load transmitted by particulates sheared away from the URT, that may manage to penetrate to the deep lungs, and more specifically, how does the viral transmission trend correlate with particulate sizes?
-This study addresses $Q_{1}$ through full-scale numerical tracking of inhaled constituents inside an anatomically accurate airway geometry with a supporting reduced-order mathematical model of the system.
-Next for $Q_{2}$, it connects the fluid dynamics findings with virological parameters, such as the sputum viral concentration [7] for a representative pathogen (SARS-CoV-2), and critiques the projected viral load transmitted to the bronchial pathways in light of the verified infectious dose of the pathogen [8, 9].
-The resulting translational analysis helps assess the proclivity for deep lung infection driven by inhalation of aerosolized intra-URT mucosal remnants.
-It is worth noting that this in silico approach, rooted in the underlying respiratory flow physics as discussed here, is agnostic to any virus specificity and is potentially extensible to other respiratory pathogens by integrating the fluid mechanics outcomes of inhaled transport with appropriate virological, epidemiological, and immunological data.
-In that context, this work could advance an emerging niche sub-field that brings together classical fluid dynamics and virology [10, 11, 12].
-For select recent relevant studies on deep lung kinetics and pathophysiology, the reader may refer to [13, 14, 15, 16, 17].
-Figure 1: (a) A demonstrative cartoon of the human respiratory system, encompassing the upper respiratory tract, the mouth, and the lower respiratory tract, extending till the deep lungs.
-The visual is adopted with copyrights’ permission from the Getty Images®.
-(b-c) Sample computed tomography (CT) imaging-based reconstruction of an adult human airway.
-It serves as a three-dimensional anatomically realistic equivalent of the cartoon in panel (a), with regions included till generation 2 (tertiary bronchi) of the tracheobronchial tree.
-The domain within the red box, i.e., panel (c), is isolated for the numerical experiments on inhaled downwind transport of microdroplets generated from the upper respiratory tract.
-The isolated region is additionally shown in panel (d) for the sagittal view with anatomical demarcations and in panel (e) for the coronal view.
+    What is the viral load transmitted to the lower airway by particulates originating from the infected intra-URT tissue surfaces?
+Could this mechanism, in general, explain the rapid onset of lung infection following the initial infection and the emergence of symptoms at URT sub-sites, such as the nasopharynx?
+First proposed (to the author’s knowledge) in [13], the veracity of such a mechanism can dispel the caveat concerning time-scale inconsistencies whereby attributing the rapid downwind progression of infection solely to tissue-level replication of the invading pathogen could be a stretch.
+This study addresses $Q_{1}$ through full-scale numerical tracking of inhaled constituents inside a representative anatomically accurate airway geometry with a supporting reduced-order anatomy-guided mathematical model of the system.
+The anatomical test geometry extends up to the third generation of bronchial branching, with cavity outlets leading into finer respiratory bronchioles; see Fig 1.
+To generate a summative assessment of bronchial deposition alongside deep lung penetration of URT-derived particulates, this study extracts and combines: (a) the simulated bronchial deposition percentages up to generation 3; and (b) the percentage of particulates progressing further downwind through the distal bronchial outlets into the deeper respiratory bronchioles.
+To note, the term deep lungs refers to the lower regions of the respiratory system where gas exchange primarily occurs.
+This domain includes the bronchioles, alveolar ducts, and alveoli [14].
+A detailed analysis of, say, the alveolar deposition trends in the deeper regions of the lungs lies beyond the scope of the present methodology.
+Using physiological modeling in a test respiratory domain, this paper exclusively assesses the lower airway penetration potential for URT-derived pathogenic microdroplets and how the mechanism could be correlated to secondary infection onset.
+However, for supplementary reading, comprehensive bronchial deposition profiles for (mostly pharmaceutical) particles administered nasally and/or orally from outside can be found in several excellent studies; e.g., see [15, 16, 17, 18, 19, 20, 21].
+Figure 1: Defining the physiological domain.
+(a) A demonstrative cartoon of the human respiratory system, encompassing the upper respiratory tract, the mouth, and the lower respiratory tract, extending till the deep lungs.
+The visual is adopted with a perpetual license agreement from the Getty Images®.
+(b) Sample computed tomography (CT) imaging-based reconstruction of an adult human airway.
+It serves as a three-dimensional anatomically realistic equivalent of the cartoon in panel (a), with regions included till generation 3 (tertiary bronchi) of the tracheobronchial tree.
+For confirmation, note that there are 8 distal outlets (see panels (b)-(d)), implying $2^{G_{n}}$ = 8, where $G_{n}=3$ is the generation number, considering two-way bifurcation for each bronchial tube at every transition [22].
+The domain within the red box is isolated for the numerical experiments on inhaled downwind transport of microdroplets generated from the upper respiratory tract.
+The isolated region is additionally shown in panel (c) for the sagittal view with anatomical demarcations and in panel (d) for the coronal view.
 The symbol $g$ signifies the gravity direction in the numerical simulations and the subsequent analytical framework, with $x$, $y$, and $z$ defining the spatial orientation of the test cavity.
-Box (c) additionally highlights the precise length scale of the geometry.
-II Methods
-II.1 Numerical modeling of inhaled transport of microdroplets generated in the upper respiratory tract
-As is the case for inhaled transport of microdroplets (also referred to as particulates, or equivalently, simply as particles in this exposition) through the respiratory cavity, in numerical simulations involving the dispersion of small particles under dilute conditions—the common approach assumes one-way coupling.
+Panel (b) additionally highlights the geometric length scale.
+Next for $Q_{2}$, the fluid dynamics findings on particulate penetration are integrated with virological parameters, such as the sputum viral concentration [23] for a representative pathogen (SARS-CoV-2), and the projected viral load transmitted to the bronchial pathways is compared against the verified infectious dose of the pathogen [24, 25].
+The resulting translational analysis helps assess the proclivity for lower airway infection that is driven by inhalation of aerosolized intra-URT mucosal remnants.
+It is worth noting that this in silico approach, rooted in the underlying respiratory flow physics as discussed here, is agnostic to any virus specificity and potentially extensible to other respiratory pathogens by integrating the fluid mechanics outcomes of inhaled transport with appropriate virological and immunological data.
+In that context, this work can contribute toward advancing an emerging niche sub-field that brings together classical fluid dynamics and virology [26, 27, 28, 29, 25, 30].
+For select recent relevant studies on deep lung kinetics and pathophysiology, the reader may refer to [31, 32, 33, 13, 34]; additionally, one may peruse [35, 36] for engaging accounts of airway flow physics, its potential impact on lung function, and the pertinent open problems.
+II Materials and methods
+Numerical modeling of inhaled downwind transport of URT-derived microdroplets
+As is the case for inhaled transport of microdroplets (also referred to as particulates, or equivalently simply at times as particles in this exposition) through the respiratory cavity, in numerical simulations involving the dispersion of small particles under dilute conditions—the common approach assumes one-way coupling.
 This reduction implies that while the airflow continuum carries the particulates, the impact of such particulates on the underlying flow regimes could be disregarded.
@@
 The simulated airflow field is coupled with the Lagrangian particle transport analysis to derive the intra-airway deposition and penetration trends.
-II.1.1 Anatomical geometry reconstruction and spatial discretization
-From de-identified, high-resolution, medical-grade computed tomography (CT) imaging, this study has first reconstructed a complete three-dimensional adult respiratory airway [18]; see Fig.
-1a-c.
+II.0.1 Anatomical geometry reconstruction, spatial discretization, and mesh sensitivity analysis
+From de-identified, high-resolution, medical-grade computed tomography (CT) imaging, this study has first reconstructed a complete three-dimensional adult respiratory airway [37]; see Fig 1a-b.
 The extraction warranted a radio-density thresholding between $-1024$ to $-300$ Hounsfield units [38, 39] to capture the airspace from the CT slices.
-Aiming to address $Q_{1}$ and $Q_{2}$ (posed in §I), the reconstruction has then been digitally redacted to focus on the space mapping the vocal fold region of the larynx (space with highest concentration of liquid particulates formed from intra-URT mucosal break-up) along with the distal laryngeal chamber, the trachea, and the lower airway extending till generation 2 of the tracheobronchial tree, with primary and tertiary bronchi, followed by the respiratory bronchiolar outlets as entry to the deeper recesses of the lungs.
-See Fig.
-1d-e for the anatomical test geometry with the described regions marked out.
-To prepare the domain for numerical simulations, the test cavity was spatially segregated into 2.5 million graded, unstructured, tetrahedral elements, along with four layers of pentahedral cells (with 0.025-mm height for each cell and an aspect ratio of 1.1) extruded at the airway cavity walls [21] to resolve the near-wall particulate dynamics; see Fig.
-2.
-Figure 2: The black box in panel (a) marks the location of the demonstrated mesh region mapping a sagittal cross-section running mid-way through the test cavity (shown henceforth in Fig.
-3).
-Panels (b) and (c) present multi-scale visuals of the mesh.
-In specific, panel (c) highlights the four layers of pentahedral cells used for near-wall refinement to ensure reliable numerical tracking of particle deposition.
-While using the LES scheme, a finer mesh refinement typically leads to more precise outcomes.
-For instance, at the upper limit, achieving results akin to Direct Numerical Simulation (DNS) is possible if the grid dimensions are smaller than the Kolmogorov scale, $\mathcal{K}$, which is defined as follows [22]:
+Aiming to address $Q_{1}$ and $Q_{2}$ (pitched in the introduction), the reconstruction has then been digitally redacted to focus on the space mapping the vocal fold region of the larynx (space with highest concentration of liquid particulates formed from intra-URT mucosal fragmentation) along with the distal laryngeal chamber, the trachea, and the lower airway extending till generation 3 of the tracheobronchial tree, incorporating the primary, secondary, and tertiary bronchi, followed by the respiratory bronchiolar outlets as entry to the deeper recesses of the lungs.
+See Fig 1c-d for the redacted test geometry with the described regions marked out.
+To prepare the domain for numerical simulations, a grid refinement analysis was conducted with the test cavity being spatially segregated into 0.5, 1.0, 1.5, 2.0, 2.5, and 3.0 million graded, unstructured, tetrahedral elements, along with four layers of pentahedral cells (with 0.025-mm height for each cell and an aspect ratio of 1.1) extruded at the airway cavity walls [40] to resolve the near-wall particulate dynamics; see Fig 2a-h.
+The variables assessed in the sensitivity study included the resistance $\mathcal{R}$ (in Pa.min/L) to the simulated inhaled airflow, calculated as $|\Delta P|/Q$, where $|\Delta P|$ in Pa represents the inlet-to-outlet pressure gradient driving the flow and $Q$ denotes the volumetric flux in L/min; the area-weighted average airflow velocity magnitude $V_{o}$ (in m/s) at the outlet surfaces of the geometry; and the transmission efficiency $\eta$ (in %) within the bronchial pathways for representative particulate sizes of 1, 5, 10, and 20 $\mathrm{\mu m}$.
+The deposition and penetration efficiency $\eta$ quantifies the cumulative bronchial transmission trend for each particulate size—summing the deposition rates at the primary, secondary, and tertiary bronchi, together with the penetration rate through the distal bronchiolar outlets into the deeper lung regions (see Fig 1c-d).
+The fluctuation trends of the tracked variables are shown in Fig 2i-k.
+Specifically, the simulation results yield:
+$\sigma(\mathcal{R})_{\rm{All}}$ $=0.0051~{}\rm{Pa.min/L},$ $\sigma(\mathcal{R})_{\rm{3}}$ $=0.0041~{}\rm{Pa.min/L},$ (1)
+$\sigma(V_{o})_{\rm{All}}$ $=0.0085~{}\rm{m/s},$ $\sigma(V_{o})_{\rm{3}}$ $=0.0026~{}\rm{m/s},$
+$\sigma(\eta)_{\rm{All}}$ $=7.94\%,$ $\sigma(\eta)_{\rm{3}}$ $=0.44\%,$
+where $\sigma(\cdot)_{\rm{All}}$ denotes the standard deviation across all six grids and $\sigma(\cdot)_{\rm{3}}$ denotes the standard deviation of the simulated data from the last 3 grids (i.e., cases with 2.0, 2.5, and 3.0 million unstructured tetrahedral cells).
+Based on the asymptotic convergence trend observed in these final three cases, the intermediate 2.5-million cell grid resolution was selected for the overall study.
+The choice is consistent with detailed studies on grid convergence and computational stability for simulating physiologically realistic airflow and particle deposition in the human respiratory system; e.g., see [41].
+Figure 2: Assessing the numerical sensitivity.
+Panel (a) marks the location of the demonstrated unstructured tetrahedral meshes spatially refined with the following cell counts (in million): (b) 0.5, (c) 1.0, (d) 1.5, (e) 2.0, (f) 2.5, (g) 3.0.
+Panel (h) highlights the four layers of pentahedral cells used for near-wall refinement.
+Panels (i)-(k) present the computed flow and particulate transport variables across the six mesh resolutions: (i) resistance $\mathcal{R}$ to inhaled airflow, (j) area-weighted average airflow velocity $V_{o}$ at the outlet faces, and (k) deposition (or, penetration) rate $\eta$ (%) in bronchial pathways for select particle sizes of 1, 5, 10, and 20 $\mu m$.
+The circle sizes to the right of panel (k) represent the number of tracked particulates in each simulation, scaled proportionally: red = 597, orange = 831, ocher = 1208, green = 1459, dark green = 1622, and blue = 1954 tracked particulates per size.
+Tracked particle counts equal the number of surface elements at inlet, directly proportional to the spatial resolution of the respective mesh.
+Notably, while using the LES scheme, a finer mesh refinement typically leads to more precise outcomes.
+For instance, at the upper limit, achieving results akin to Direct Numerical Simulation (DNS) is possible if the grid dimensions are smaller than the Kolmogorov scale, $\mathcal{K}$, defined as [42]:
 $\mathcal{K}=\left(\frac{\nu^{3}}{\varepsilon}\right)^{1/4},$ (2)
 with $\varepsilon$ as the turbulence dissipation rate and $\nu$ being the fluid kinematic viscosity.
-Another length scale of note is the Taylor scale, $\lambda$, which typically exceeds $\mathcal{K}$ and is defined as
+Another length scale of note is the Taylor scale, $\lambda$, which typically exceeds $\mathcal{K}$ and is defined as:
 $\lambda=\left(\frac{10\nu k}{\varepsilon}\right)^{1/2},$ (3)
 with $k$ being the turbulence kinetic energy.
-From the simulation data of the inhaled airflow field from this study, it is seen that both $\lambda$ and $\mathcal{K}$ collapse to $\mathcal{O}(10^{-4})$ m, while the mean grid scale is also $\rightarrow\mathcal{O}(10^{-4})$ m, suggesting that the test grid has been sufficiently resolved for precise estimation of the transport parameters.
-II.1.2 Numerical simulation of inhaled airflow
-We have employed the LES approach to numerically model the inhaled airflow, with eddies exceeding the grid scale explicitly resolved, whereas those that are smaller than the grid scale are approximated.
+From the simulation data of the inhaled airflow field, it is seen that both $\lambda$ and $\mathcal{K}$ collapse to $\mathcal{O}(10^{-4})$ m, while the mean grid scale is also $\rightarrow\mathcal{O}(10^{-4})$ m, suggesting that the test grid has been sufficiently resolved for reliable estimation of the transport parameters.
+II.0.2 Numerical simulation of inhaled airflow
+We have implemented the LES approach to numerically model the inhaled airflow, with eddies exceeding the grid scale explicitly resolved, whereas those that are smaller than the grid scale are approximated.
 Specifically, fluctuations below the grid size, referred to as subgrid scales, are filtered out, and their impact on larger scales is replicated through modeling.
-Assuming incompressible and isothermal conditions for the inhaled air flux, the filtered continuity and Navier–Stokes equations are respectively as follows:
+Assuming incompressible, isothermal conditions for the inhaled air flux, the filtered continuity and Navier–Stokes equations are respectively:
 $\frac{\partial}{\partial x_{i}}\left(\rho\overline{u}_{i}\right)=0$ (4)
@@
 with the model constants in the previous equations, i.e., $C_{k}$ and $C_{\varepsilon}$, being determined dynamically [43].
-Applying the LES scheme as described above, the inhaled airflow through the anatomical airspace was replicated for 15 L/min inhalation rate [27].
-See Fig.
-1d for the pressure inlet face at the location of vocal folds in the larynx and for the pressure outlets at the distal ends of the reconstructed generation 2 bronchial tubes.
-Enforcing no slip (i.e., zero velocity) boundary condition at the airway walls, the pressure gradient-driven simulation used time-steps of 0.0002 s [25], which is smaller than the Kolmogorov time scale = $\left(\nu/\varepsilon\right)^{1/2}$ [22], for a flow solution time of 0.35 s.
+Applying the LES scheme as described above, the inhaled airflow through the anatomical airspace was replicated for 15 L/min inhalation rate, conforming with normal relaxed breathing conditions [47].
+See Fig 1c for the pressure inlet face at the location of vocal folds in the larynx and for the pressure outlets at the distal ends of the reconstructed generation 3 bronchial tubes.
+Enforcing no slip (i.e., zero velocity) boundary condition at the airway walls, the pressure gradient-driven simulation used time-steps of 0.0002 s.
+The latter was chosen based on reported findings [45] on the time-step size warranted to fully resolve the unsteady turbulent airflow field in a realistic upper airway model, it being smaller than the Kolmogorov time scale = $\left(\nu/\varepsilon\right)^{1/2}$ [42], for a flow solution time of 0.35 s.
 The simulation was executed on a segregated solver with pressure-velocity coupling and second-order upwind spatial discretization.
+For the transient formulation, a bounded second-order implicit scheme was employed, to strike a balance between accuracy (due to the second-order formulation), stability (from the implicit approach), and boundedness (to prevent non-physical oscillations).
 The eventual solution convergence was monitored by minimizing the mass continuity residual to $\mathcal{O}(10^{-3})$ and the velocity component residuals to $\mathcal{O}(10^{-6})$.
-Also, considering the warmed-up state of inhaled air passing through the respiratory pathway, the air density $\rho$ was set at 1.204 kg/m3 in the simulations, with 18.218$\times$10-6 m2/s as its kinematic viscosity $\nu$.
-II.1.3 Numerical experiments for particulate transport
-Against the solved airflow field, the particulates assumed to be formed from intra-URT mucus break-up were tracked from the vocal fold region of the larynx (see the marked inlet face in Fig.
-1d).
-The air-particle phases were one-way coupled with the particles being impacted by the ambient flow field; the underlying flow domain was considered quasi-steady while evaluating the particle transport parameters.
+Also, considering the warmed-up state of inhaled air passing through the respiratory pathway, the air density $\rho$ was set at 1.204 kg/m3 in the simulations, with 15.16$\times$10-6 m2/s as its kinematic viscosity $\nu$.
+II.0.3 Numerical experiments for particulate transport
+Against the solved airflow field, the particulates formed intra-URT were tracked from the vocal fold region of the larynx (see the marked inlet face in Fig 1c).
+The air-particle phases were one-way coupled with the particles (assumed spherical) being impacted by the ambient flow field; the underlying flow domain was considered quasi-steady while evaluating the particle transport parameters.
 Lagrangian-based inert discrete phase model, with a Runge-Kutta solver, was used to numerically integrate the particle transport equation:
 $\frac{du_{pi}}{dt}=\frac{18\mu}{d^{2}\rho_{p}}\frac{C_{D}Re_{p}}{24}\left(u_{i}-u_{pi}\right)+g_{i}\left(1-\frac{\rho}{\rho_{p}}\right)+F_{i}.$ (11)
-Here $u_{pi}$ represents the particulate velocity, $\rho_{p}$ is the material density of the particulates, $d$ represents the particulate diameter, $Re_{p}$ is the particulate Reynolds number, $g_{i}$ signifies the gravitational acceleration in the $i$ direction, $C_{D}$ is the drag coefficient, and $F_{i}$ is any other additional body forces per unit particulate mass, e.g., the Saffman lift force exerted by a typical flow-shear field on small particulates transverse to the airflow direction.
+Here $u_{pi}$ represents the particulate velocity, $\rho_{p}$ is the material density of the particulates, $d$ represents the particulate diameter, $Re_{p}$ is the particulate Reynolds number, $g_{i}$ signifies the gravitational acceleration in the $i$ direction, $C_{D}$ is the drag coefficient, and $F_{i}$ represents additional body forces per unit particulate mass in the form of the Saffman lift force exerted by a typical flow-shear field on small particulates transverse to the airflow direction.
 In this context, note that the solution scheme considered the particulates to be large enough to ignore any Brownian motion effects on their dynamics.
-To evaluate the drag term (i.e., the co-factor on the first term) on the right hand side of equation 10, the quantities $Re_{p}$ and $C_{D}$ are respectively computed as
-$Re_{p}$ $=\frac{\rho_{p}d\left|u_{i}-u_{pi}\right|}{\mu}$ $C_{D}$ $=a_{1}+\frac{a_{2}}{Re_{p}}+\frac{a_{3}}{{Re_{p}}^{2}},$ (11)
-where $\mu$ is the molecular viscosity of the ambient fluid (i.e., air), while $a_{1}$, $a_{2}$, and $a_{3}$ are functions of $Re_{p}$ determined based on the spherical drag law [28].
+To evaluate the drag component in equation 11, the quantities $Re_{p}$ and $C_{D}$ are respectively computed as: $Re_{p}=\rho_{p}d\left|u_{i}-u_{pi}\right|/\mu$ and $C_{D}=a_{1}+a_{2}/Re_{p}+a_{3}/{Re_{p}}^{2}$; where $\mu$ is the molecular viscosity of the ambient fluid (i.e., air), while $a_{1}$, $a_{2}$, and $a_{3}$ are functions of $Re_{p}$ determined based on the spherical drag law [48].
 Subsequently, the particulate trajectories are derived from their spatiotemporal locations, $x_{i}\left(t\right)$, obtained by numerical integration of the following velocity vector equation:
@@
 The numerical experiments in this study tested particulates of diameters $\mathrm{1-30~{}\mu m}$ (with increments of 1 $\mathrm{\mu m}$).
-In total, $\mathcal{N}$ = 1622 particulates of each size were tracked; $\mathcal{N}$ being the number of surface elements in the mesh layer mapping the inlet face (see Fig.
-1).
+In total, $\mathcal{N}$ = 1622 particulates of each size were tracked; $\mathcal{N}$ being the number of surface elements in the mesh layer mapping the inlet face (see Fig 1).
 Also, considering that saliva-mixed mucus is 99.5% water [49], the material density of the particulates was assumed to be $\rho_{p}$ = 1.00175 g/ml, approximated as the weighted average (see equation 13 below) between 99.5% water and the residual pathogenic and biological non-volatile compounds with a representative density of $\rho_{nv}=$ 1.35 g/ml; e.g., for protein [50], the density being independent of the nature of the protein and particularly independent of its molecular weight.
@@
 The modeling approach also discounted any heat transfer effects between the flow constituents and the surrounding tissues enclosing the anatomical airspace.
-Figure 3: (a) Laryngotracheal region upwind from the primary bronchi used to develop a simulation-informed reduced-order analytical model (S-ROAM) for bronchial transmission.
-AA marks the distal larynx upwind face and BB marks the downwind face of the distal tracheal cavity; see labels in Fig.
+II.0.4 Supplementary ‘diversion’: bronchial transmission trend for particulates inhaled from outside
+While the focus of this study is singularly on the downwind bronchial transmission of particulates generated within the URT, it would be of scholarly interest to undertake a brief detour at this point to explore an auxiliary question (that has been nonetheless well-explored in literature): for the particulates inhaled from the external environment, which particulate sizes would be efficient at penetrating to the bronchial spaces?
+Here, we address this question specifically for the test geometry used in this study, through transport simulations within its non-redacted complete anatomy (Fig 1b) wherein inhaled particulates are entering the airway through both the left and right nostrils.
+Comparing the cavity spaces in the complete airway geometry (Fig 1b) and the redacted model (Fig 1c-d), the volume of the former is approximately 3.48 times larger than that of the redacted system.
+Accordingly, following the described meshing protocol, the full airway was discretized into 2.5 million $\times$ 3.48 = 8.7 million unstructured tetrahedral elements, with four layers of pentahedral cells lining the cavity surfaces.
+Using this meshed geometry, particulate transport was simulated at an inhaled airflow rate of 15 L/min, following methods outlined earlier.
+The only modification was setting the material density of the inhaled particulates to 1.3 g/ml, consistent with typical values for environmentally dehydrated respiratory ejecta [49, 25], which constitutionally form the pathogen-bearing particulates inhaled from external air.
+Fig 3 demonstrates the bronchial transmission trend for particulates inhaled from outside.
+Therein, panel (a) shows the cumulative transmission percentage, $\eta$, at the primary, secondary, and tertiary bronchi, together with the penetration rate (in %) through the geometry outlets that serve as entry to the respiratory bronchioles.
+Panel (b) solely presents $\eta_{d}$, the penetration rate through the bronchiolar outlets, indicating the proportion of inhaled particulates reaching the deeper lung regions.
+Considering select particulate sizes, $\eta$ starts at 99.36% for particulates of diameter $d=1~{}\mathrm{\mu m}$, decreases to 96.86% for $d=5~{}\mathrm{\mu m}$, to 31.46% for $d=10~{}\mathrm{\mu m}$, and approaches 0% for $d=15~{}\mathrm{\mu m}$ and larger.
+Remarkably though, as we will find later (in the results), for particulates generated within the URT, the quantity $\eta$ shoots up significantly, approaching 90% for $d=10~{}\mathrm{\mu m}$ and steadily sustaining at $>$ 60% even for $d=15~{}\mathrm{\mu m}$.
+Figure 3: Transmission trend for particulates inhaled from outside and navigating the complete anterior airway.
+(a) $\eta$ represents the cumulative deposition percentages at the primary, secondary, and tertiary bronchi, together with the penetration rate through the bronchiolar outlets into the deeper lung regions; (b) $\eta_{d}$ isolates the penetration rate through the bronchiolar outlets with the tracked particulates moving into the deeper lung regions.
+The reader may find it insightful to compare the trend reported here with Fig 6a-b on the navigation trend for particulates generated within the URT (and still airborne at the larynx).
+In the latter scenario, the larger particulates exhibit much greater efficiency at penetrating to the lower airway.
+II.1 A simplified analytical model for inhaled transport through the laryngotracheal space
+To verify the particulate transport trends to the bronchial domain as derived numerically, let us invoke a simulation-informed reduced-order analytical model (S-ROAM), with a two-dimensional channel mimicking the laryngotracheal domain.
+Therein, the inhaled airflow is modeled in the dimensionless complex $\chi$ plane, with $\chi=\alpha+\mathrm{i}\,\beta$ and $\mathrm{i}\,\beta$ aligned with the channel’s streamwise axis; see Fig 4.
+The two-dimensional channel has its streamwise length and cross-stream width based on the averaged dimensions of the anatomical tract from Fig 4a-b.
+The channel walls are placed at $\alpha=0,1$.
+Based on the geometric inputs, the S-ROAM enforces $L/W\approx 9$, where $L$ represents the streamwise channel length and $W$ is the inlet width.
+For a detailed exposition of this classical modeling approach for complex respiratory systems, see our recent preprint [51].
+Figure 4: Defining the analytical domain, with representative numerical visuals.
+(a) Laryngotracheal region upwind from the primary bronchi used to develop a simulation-informed reduced-order analytical model (S-ROAM) for bronchial transmission.
+AA marks the larynx upwind face and BB marks the downwind face of the distal tracheal cavity; see labels in Fig.
 1d.
@@
 The linear spatial distance between the two faces is approximately 115.84 mm, at angle $\theta_{1}$ = 23.77 to the vertical (direction of gravity in the numerical simulations).
-Panel (b) highlights the location and orientation of the planar cross-section shown in panel (c) and also previously as the location of the mesh visuals in Fig.
-2.
+Panel (b) highlights the location and orientation of the planar cross-section shown in panel (c) and also previously as the location of the mesh visuals in Fig 2.
 The vortex strengths and positions for the S-ROAM are extracted from the simulated data mapping this two-dimensional plane; the plane cuts through the entire cavity of panel (a).
-$\theta_{2}$ = 33.94 is the assumed angle in S-ROAM, at which microdroplets enter the AA inlet face and is governed by the anatomical shape of the cavity upwind from AA (marked by the dashed black traces; also see Fig.
-1).
-II.2 Analytical model for inhaled airflow and particulate transport
-To verify the particulate transport trends to the bronchial domain as derived numerically, this study invokes a simulation-informed reduced-order analytical model (S-ROAM), with a two-dimensional channel mimicking the laryngotracheal domain.
-Therein, the inhaled airflow was modeled in the dimensionless complex $\chi$ plane, with $\chi=\alpha+\mathrm{i}\,\beta$ and $\mathrm{i}\,\beta$ aligned with the channel’s streamwise axis; see Fig.
-3.
-The two-dimensional channel had its streamwise length and cross-stream width based on the averaged dimensions of the anatomical tract from Fig.
-3a-b.
-The channel walls were placed at $\alpha=0,1$.
-Based on the geometric inputs, the S-ROAM enforced $L/W\approx 9$, where $L$ represents the streamwise channel length and $W$ is the inlet width.
-For additional details on the analytical approach, see our recent paper [31].
-The reduced system modeled an extended vortex patch emerging in the full-scale numerical simulation as a collection of five point vortices distributed over the area of the simulated vortex core (on the channel plane, mapping the mid-sagittal section of the three-dimensional cavity; see Fig.
-3b), bearing dimensionless circulations $\Gamma_{i}=\omega A/5UW$, where $\omega$ was the mean vorticity (as determined from the numerical simulations) spread over the patch area $A$ and $U$ is the characteristic streamwise air speed through the channel.
-From the simulated data, we had: $\omega=1750$ s-1, $A=2.65\times 10^{-5}$ m2, and $U=Q/a=2.16$ m/s, with $a=$ 115.82 mm2 being the area of the inlet face and $Q=15$ L/min $=0.00025$ m3/s being the simulated inhalation flux.
-The anatomical planar section from which the vorticity information was extracted is shown in Fig.
-3c.
+$\theta_{2}$ = 33.94 is the assumed angle in S-ROAM, at which microdroplets enter the AA inlet face and is governed by the anatomical shape of the cavity upwind from AA (marked by the dashed black traces; also see Fig 1).
+(d) 30 randomly selected representative velocity streamlines extracted from the numerically simulated inhaled airflow field.
+(e) Simulated trajectory of a representative 10-$\mathrm{\mu m}$ particulate.
+The geometric centroid of the inlet face, marked by the solid circle, was chosen as the particulate’s position at tracking time $t=0$.
+The reduced system modeled the core of an extended vortex patch emerging in the full-scale numerical simulation as a random collection of five point vortices distributed over the area of the simulated vortex core (on the channel plane, mapping the mid-sagittal section of the three-dimensional cavity; see Fig 4b), bearing dimensionless circulations $\Gamma_{i}=\omega A/5UW$, where $\omega$ was the mean vorticity (as determined from the numerical simulations) spread over the patch area $A$ and $U$ is the characteristic streamwise air speed through the channel.
+From the simulated data, we had: $\omega\approx 1750$ s-1, $A\approx 2.65\times 10^{-5}$ m2, and $U=Q/a=2.16$ m/s, with $a=$ 115.82 mm2 being the area of the inlet face and $Q=15$ L/min $=0.00025$ m3/s being the simulated inhalation flux.
+Based on the angular orientation of the modeled core in the simulated vorticity field, the model point vortices were placed on a line that subtended $\approx$ 2.2 in counter-clockwise sense with the vertically downward direction.
+The anatomical planar section from which the vorticity information was extracted is shown in Fig 4c.
 In the S-ROAM, the straight streamlines at the channel walls were established through inclusion of periodic images of the point vortices in the cross-stream $\alpha$ direction.
@@
 with $\phi$ as the velocity potential and $\psi$ as the real-valued flow streamfunction; the asterisk denotes complex conjugation.
-Subsequently, the inhaled particulate motion was analytically derived using a simplified version of the Maxey-Riley equation [35], in the two-dimensional vector form [36, 37]:
+Subsequently, the inhaled particulate motion was analytically derived using a simplified version of the Maxey-Riley equation [55, 56], in the two-dimensional vector form [57, 58]:
 $\frac{d{\mbox{$w$}}}{dt}=-\left[{\mbox{$J$}}+\frac{2\,{St}^{-1}}{3(\sigma+1)}{\mbox{$I$}}\right]\cdot{\mbox{$w$}}+\frac{\sigma}{\sigma+1}\left({Fr}^{-2}{\mbox{$g$}}-\frac{\text{D}{\mbox{$u$}}}{\text{D}t}\right).$ (15)
@@
 This form of the model ignores the Faxen correction terms and the Basset-Boussinesq history force.
-Particulates are assumed to be entering the S-ROAM domain (inclined at an angle $\theta_{1}=23.77^{\circ}$ to the direction of gravity, per the general shape and orientation of the test anatomical cavity) with speed $U$ and at an angle of $\theta_{2}=33.94^{\circ}$ with respect to the S-ROAM’s $\alpha$-axis, mimicking the simulated motion of particles entering the region; see Fig.
-3c.
-As representative examples, particulate trajectories were derived for two initial positions at the mid-point and near the right edge of the reduced channel’s inlet span, and for seven different particulate diameters, namely, 5, 10, 15, 20, 25, 30, and (as an extremal case) 50 $\mathrm{\mu m}$.
-II.3 Connecting the fluid dynamics outcomes to virology
-From the numerical experiments, this study has deduced the simulated deposition and penetration efficiencies of the particulates along the bronchial tubes (see Table 1) and evaluated the respective volumetric transmission to the bronchi.
-The projected net deposited (at the primary and tertiary bronchi) and penetrated (moving into the respiratory bronchioles) volumes are then multiplied with the sputum viral concentration for a specific pathogen (in this study, SARS-CoV-2) to evaluate and compare the viral load transported via inhaled aerial advection of URT-derived liquid particulates to the lower airway and deep lungs, for select test particulate sizes, namely, 1, 5, 10, and 15 $\mathrm{\mu m}$.
-Figure 4: 30 representative velocity streamlines extracted from the numerically simulated inhaled airflow field.
-Note that the average sputum viral concentration for SARS-CoV-2 has been reported as $\mathcal{V}$ = 7.0$\times$106 virions/ml, through count measurements of the RNA copies for the single-stranded virus present in the airway liquid samples collected from hospitalized COVID-19 patients [7].
+Particulates are assumed to be entering the S-ROAM domain (inclined at an angle $\theta_{1}=23.77^{\circ}$ to the direction of gravity, per the general shape and orientation of the test anatomical cavity) with speed $U$ and at an angle of $\theta_{2}=33.94^{\circ}$ with respect to the S-ROAM’s $\alpha$-axis, mimicking the simulated motion of particles entering the region; see Fig 4c.
+As representative examples, particulate trajectories were derived for two initial positions at the mid-point and near the right edge of the reduced channel’s inlet span, and for seven different particulate diameters, namely 5, 10, 15, 20, 25, 30, and (as an extremal case) 50 $\mathrm{\mu m}$.
+II.2 Connecting the fluid dynamics outcomes to pathogen-specific virological data
+From the numerical experiments, this study has deduced the simulated deposition and penetration efficiencies of the URT-derived particulates along the bronchial tubes (see Table 1) and evaluated the respective volumetric transmission to the bronchi.
+The projected net deposited (at the primary, secondary, and tertiary bronchi) and penetrated (moving into the respiratory bronchioles) volumes are then multiplied with the sputum viral concentration for a specific pathogen (in this study, SARS-CoV-2) to evaluate and compare the viral load transported via inhaled aerial advection of URT particulates to the lower airway and deep lungs, for select test particulate sizes, namely, 1, 5, 10, and 15 $\mathrm{\mu m}$.
+Note that the average sputum viral concentration for SARS-CoV-2 has been reported as $\mathcal{V}$ = 7.0$\times$106 virions/ml, through count measurements of the RNA copies for the single-stranded virus present in the airway liquid samples collected from hospitalized COVID-19 patients; see [23].
 III Results
-Table 1: Numerically simulated bronchial deposition and penetration data: The particulate sizes, that have been explored further in Table 2 (to formalize the viral transmission trends as a function of the microdroplet dimensions), are in bold font.
-Symbols: $d$ = tested aerosol (or, droplet) diameter; $\mathcal{N}$ = total number of aerosols/droplets tracked for each particulate diameter; $n_{p}$ = number of deposited particulates in the primary bronchi; $n_{t}$ = number of deposited particulates in the tertiary bronchi; $n_{d}$ = number of particulates penetrating into the respiratory bronchioles toward the deep lungs; $\eta$ = deposition (or, penetration) rate (in %) to the bronchial pathways.
-Separated aerosol/droplet size
+Table 1: Numerical transmission trend: The numerically simulated bronchial deposition and penetration data are detailed herein.
+The particulate sizes, that have been explored further in Table 2 (to formalize the viral transmission trends as a function of the microdroplet dimensions), are in bold font.
+Symbols: $d$ = tested aerosol (or, droplet) diameter; $\mathcal{N}$ = total number of aerosols (or, droplets) tracked for each particulate diameter; $n_{p}$ = number of deposited particulates in the primary and secondary bronchi; $n_{t}$ = number of deposited particulates in the tertiary bronchi; $n_{d}$ = number of particulates penetrating into the respiratory bronchioles toward the deep lungs; $\eta$ = cumulative deposition (or, penetration) rate (in %) to the bronchial pathways and computed as $100\times\left(\sum_{j}n_{j}\right)/\mathcal{N}$, with $j\in\{p,t,d\}$.
+Separated particulate size
 ($d$, in $\mathrm{\mu m}$)
@@
 $n_{d}$
-Lung penetration ($\eta$, in %)
+Bronchial
+transmission
+($\eta$, in %)
+1 1622 6 0 1616 100.00
+2 1622 33 3 1586 100.00
 3 1622 11 0 1611 100.00
-2 1622 4 0 1617 99.94
-3 1622 14 0 1608 100.00
-4 1622 16 0 1606 100.00
-5 1622 32 3 1587 100.00
-6 1622 83 13 1526 100.00
-7 1622 123 25 1469 99.69
-8 1622 170 41 1381 98.15
-9 1622 173 34 1291 92.36
-10 1622 210 41 1183 88.41
-11 1622 269 47 1054 84.46
-12 1622 316 43 933 79.65
-13 1622 369 47 797 74.78
-14 1622 420 88 607 68.74
-15 1622 461 93 433 60.85
-16 1622 489 89 293 53.70
-17 1622 460 91 217 47.35
-18 1622 417 99 144 40.69
-19 1622 365 91 80 33.05
-20 1622 324 78 37 27.07
-21 1622 280 59 17 21.95
-22 1622 221 48 12 17.32
-23 1622 172 17 10 12.27
-24 1622 120 4 1 7.71
-25 1622 81 3 0 5.18
-26 1622 57 1 0 3.58
+4 1622 14 1 1607 100.00
+5 1622 33 3 1586 100.00
+6 1622 53 10 1559 100.00
+7 1622 91 34 1493 99.75
+8 1622 136 34 1422 98.15
+9 1622 168 51 1277 92.23
+10 1622 214 39 1175 88.04
+11 1622 229 55 1083 84.28
+12 1622 300 56 930 79.28
+13 1622 368 59 787 74.85
+14 1622 424 60 634 68.93
+15 1622 449 93 453 61.34
+16 1622 445 90 321 52.77
+17 1622 433 111 217 46.92
+18 1622 395 108 146 40.01
+19 1622 361 104 72 33.11
+20 1622 309 86 45 27.13
+21 1622 268 63 25 21.95
+22 1622 218 49 13 17.26
+23 1622 168 22 10 12.33
+24 1622 118 4 1 7.58
+25 1622 83 0 0 5.12
+26 1622 58 0 0 3.58
 27 1622 38 0 0 2.34
-28 1622 20 0 0 1.23
-29 1622 5 0 0 0.31
+28 1622 19 0 0 1.17
+29 1622 3 0 0 0.18
 30 1622 0 0 0 0.00
-III.1 Bronchial deposition and deep lung penetration as projected from numerical modeling
-Figure 4 presents sample inhaled airflow velocity streamlines from the numerical simulation of inhaled airflow passing through the CT-reconstructed anatomical cavity.
-The simulated inhalation rate of 15 L/min to mimic relaxed breathing warranted an inlet-to-outlet pressure gradient of -9.5 Pa.
-Against the flow field, the test particulates were tracked with their starting locations uniformly distributed on the cross-sectional space spanning the vocal fold region of the laryngeal cavity (commensurate with their hypothesized formation sites at the URT—through break-up of mucus strata along the nasopharynx, the oropharynx, and the vocal folds).
-Table 1 details the lower airway deposition and penetration data for the tested particulates bearing diameters $\mathrm{1-30~{}\mu m}$ (with increments of $\mathrm{1~{}\mu m}$).
-While it is expected that the particulates $\lesssim$ $\mathrm{5~{}\mu m}$ would relatively comfortably penetrate to the deep lungs (as is clearly the case per Table 1), the high transmission percentages, e.g., for even the 10- and 15-$\mathrm{\mu m}$ particulates is striking—them being 88.41% and 60.85%, respectively.
-In the table (for 1622 particulates tracked for each size), $n_{p}$ quantifies the number of particulates depositing in the primary bronchi, $n_{t}$ quantifies the number of particulates depositing at the tertiary bronchi, and $n_{d}$ measures the number of particulates escaping through the respiratory bronchiolar outlets of the in silico model and subsequently moving into the deeper recesses of the lungs.
-See Fig.
-1 to review the anatomical zone nomenclature for this study and the respective locations of the different physiological regions along the airway.
+III.1 Bronchial deposition and lower airway penetration as projected from numerical modeling
+Fig 4d presents sample inhaled airflow velocity streamlines from the numerical simulation of inhaled airflow.
+The simulated flux of 15 L/min warranted an inlet-to-outlet pressure gradient of -9.5 Pa.
+Against the flow field, the test particulates were tracked with their starting locations uniformly distributed on the cross-sectional space spanning the vocal fold region of the laryngeal cavity (commensurate with their hypothesized formation sites at the URT—through breakup of mucus strata along the nasopharynx, the oropharynx, and the vocal folds).
+A representative particulate trajectory for $d=10~{}\mathrm{\mu m}$ has been additionally shown in Fig 4e.
+At tracking time $t=0$, the particulate was assumed to be positioned at the geometric centroid of the inlet face.
+Eventually, the sample particulate penetrates through the outlet to respiratory bronchioles, thereby moving into the deeper lung recesses.
+Fig 5a-c and Table 1 detail the lower airway deposition and penetration data for the tested particulates bearing diameters $\mathrm{1-30~{}\mu m}$ (with increments of $\mathrm{1~{}\mu m}$).
+While it is expected that the particulates $\lesssim$ $\mathrm{5~{}\mu m}$ would comfortably penetrate to the deeper regions of lungs (as is clearly the case per Table 1; see the top rows), the high transmission percentages, e.g., for even the 10- and 15-$\mathrm{\mu m}$ particulates is striking—them being 88.04% and 61.34%, respectively.
+Figure 5: Numerically simulated transmission trend for URT-derived particulates navigating through the redacted test geometry.
+(a) $\eta$ represents the cumulative deposition percentages at the primary, secondary, and tertiary bronchi, as well as the penetration rate through the bronchiolar outlets into the deeper lung regions; (b) $\eta_{d}$ isolates the penetration rate through the bronchiolar outlets; (c) The solid red line is a fitted curve of the Heaviside step function form (see equation 17), with the solid black circles showing simulation-derived $\eta$ values (reported in Table 1).
 III.2 Viral load transmitted to the lower airway
 Table 2 lists the viral load transmitted to the bronchial pathways, for the representative test particulate sizes 1, 5, 10, and 15 $\mathrm{\mu m}$.
-For the time scale $T$ (in hours) over which the viral load transmission is to be estimated, we have used 3 days (i.e., 72 hours) based on reported data [38] on the typical time interval that has revealed confirmed infection onset in the deep lungs subsequent to the emergence of initial symptoms along the URT.
-Further, applying $\delta t$ = 5 s as the average duration for a complete breathing cycle [39], the viral load, $L_{v}$, transmitted to the bronchial pathways and deep lungs can be computed as:
+For the time scale $T$ (in days) over which the viral load transmission is to be estimated, we have used 3 days based on reported data [59] on the typical time interval that has revealed confirmed infection onset in the deep lungs subsequent to the emergence of initial symptoms along the URT.
+Further, applying $\delta t$ = 5 s as the average duration for a complete breathing cycle [60], the viral load, $L_{v}$, transmitted downwind can be computed as:
 $L_{v}=\frac{144\,\pi\,d^{3}\,\eta\,T\,\mathcal{V}\,\mathcal{P}\left(d_{i}\right)}{\delta t}\times 10^{-12}.$ (16)
@@
 For simplicity, the data in Table 2 enforced $\mathcal{P}$ = 1, irrespective of the particulate size, i.e., the particulate formation rate was conservatively assumed to be 1 per breathing cycle, for each test particulate size.
-Finally, note that while the reported numbers in Table 2 are based on the estimates guided by equation 16, the zero viral load assessment for the 1-$\mathrm{\mu m}$ particulates implies that the corresponding transmission estimate from equation 16 resulted in a fractional number $\ll$ 1, and bears no physical significance.
+Also, while the reported numbers in Table 2 are based on the estimates guided by equation 16, the zero viral load assessment for the 1-$\mathrm{\mu m}$ particulates implies that the corresponding transmission estimate from equation 16 resulted in a fractional number $\ll$ 1, and bears no physical significance.
+To note here additionally, for the assumed $\mathcal{P}$, the quantity $L_{v}$ exceeded 300 for $d\in[13,22]~{}\mathrm{\mu m}$.
+See the conclusion section later for the related translational relevance in infection mechanics.
 Table 2: Magnitude of viral load transmission into bronchial pathways as a function of inhaled particulate sizes: The duration in consideration is 3 days, with a conservative estimate of 1 particulate of each test size generated during each breathing cycle.
 The breathing cycles last 5 s [60].
-Separated aerosol/droplet size
+Separated particulate size
 ($d$, in $\mathrm{\mu m}$)
-Number of virions ferried to the bronchial pathways
+Number of virions ferried
+to the bronchial pathways
 ($L_{v}$)
-15 390
-10 168
+15 394
+10 167
 5 24
 1 0
-Figure 5: Log-log plot for $\eta$, summing up the bronchial deposition and deep lung penetration percentages (see Table 1), for each numerically tracked microdroplet size, $d$.
-The simulation-derived data points are marked by the dark blue circles.
-The solid black line is a fitted curve of the Heaviside step function form (see equation 17).
 III.3 Heaviside trend governing bronchial deposition and penetration
-Collapsing the net particulate transmission efficiencies from Table 1, labeled as $\eta$ (in %), on to a log-log scale with particulate diameters $d$ (in $\mathrm{\mu m}$) placed along the horizontal axis, reveals (as is expected) a high lower airway penetration for smaller particulate sizes, followed by a gently sloping dip; see Fig.
-5.
-The behavior could be approximately mathematized with a modified Heaviside function [40] of the following analytical form:
-$\eta(d)=\mathcal{C}_{1}+e^{\mathcal{C}_{2}/\left[1+e^{2k\left(d+\mathcal{C}_{3}\right)}\right]},$ (17)
+Collapsing the transmission efficiencies from Table 1 and Fig.
+5a on a 2D plane with particulate diameters $d$ (in $\mathrm{\mu m}$) along the horizontal axis reveals an S-trend with (as expected) a high bronchial penetration for smaller particulate sizes, followed by a gently sloped dip; see Fig 5c.
+The behavior could be approximated with a modified Heaviside function [61] of the following empirical form:
+$\eta(d)=\mathcal{C}_{1}+e^{\mathcal{C}_{2}/\left[1+e^{k\left(d+\mathcal{C}_{3}\right)}\right]},$ (17)
 where $k$ and $\mathcal{C}_{i}$, with $i\in\{1,2,3\}$, are constant fitting parameters.
-For the solid black curve plotted in Fig.
-5 and fitting through the data-points from the numerical experiments, the following values were used:
-$\mathcal{C}_{1}=-1.95,\quad\mathcal{C}_{2}=4.95,\quad\mathcal{C}_{3}=-25.0,\quad k=0.08.$ (18)
-Figure 6: (a) Simulated vorticity contour map on a representative two-dimensional cross-section (the S-ROAM domain) running approximately midway through the cavity; see Fig.
-3b.
-The region with the dominant vortex core is marked by $\mathbb{V}$.
-(b-h) show two sample particle pathlines (in red) against the S-ROAM streamlines (in grey), respectively for particle sizes 5, 10, 15, 20, 25, 30, and 50 $\mathrm{\mu m}$.
+The adopted multivariate Heaviside formalism is particularly suitable for modeling the inverted S-shaped trend of bronchial deposition and penetration as a function of the particulate sizes, owing to its ability to represent sharp yet smooth transitions between high and low deposition rates across different size thresholds.
+This function effectively captures the nonlinear behavior implicit to respiratory systems, where smaller particulates penetrate deeply into the bronchial pathways, while larger ones tend to deposit more proximally or are filtered out.
+The mathematical ansatz allows for adjustable steepness and transition points, rendering it flexible enough to fit the specific inflection points associated with particle size changes.
+Herein, the solid red curve, fitting through the numerical data-points, was derived using the Nelder-Mead simplex algorithm for systematic error minimization [62, 63].
+The algorithm, also known as the downhill simplex method, is an iterative, heuristic optimization technique designed for unconstrained optimization problems, especially where derivative information is not available or the function is noisy or discontinuous.
+The curve in Fig 5c invokes the following fitting parameter values:
+$\mathcal{C}_{1}=-2.1706,\quad\mathcal{C}_{2}=4.6755,\quad\mathcal{C}_{3}=-23.9592,\quad k=0.23468.$ (18)
+III.4 Analytical projections for particulate transport: consistent with the full-scale numerical findings
+Fig 6a demonstrates the vorticity field mapped over the section shown in Fig 4c.
+In the reduced-order analytical setup, the vortex patch core marked by $\mathbb{V}$ is modeled with a spatial assembly of five point vortices, as described in the methods.
+In Fig 6b-h, the grey curves show the streamlines of the background flow field in the S-ROAM, with the red curves tracing the sample particulate trajectories.
+Smaller particulates, owing to their low inertia and smaller $St$, follow the streamlines (on which they were embedded at entry points) more closely.
+For larger particulates, inertia-dominated dynamics, combined with gravitational impaction, tend to inhibit downwind penetration by biasing their motion toward deposition on the walls (see panels g and especially h, in Fig 6).
+However, some sufficiently large particles, such as those with diameters of 10 and 15 $\mathrm{\mu m}$ (see Fig 6c-d), can still effectively maneuver around the vortex trap to reach the lower airspace.
+This trend aligns with the full-scale numerical projections.
+Note that the particulates entering through the middle of the inlet face of the S-ROAM are marked as $\mathbb{I}$ while those entering near the right edge are marked as $\mathbb{II}$.
+As shown in Fig 7, the particulates $\mathbb{II}$, being nearer to the vortex region, are deviated more (from the airflow streamlines they were embedded on at the channel inlet) compared to the particulates $\mathbb{I}$.
+The deviation generically grows as the particulate diameters are increased, which aligns with the numerical findings reported in Table 1.
+A greater deviation implies that the particulates are being increasingly shifted toward the channel walls, which in turn would result in declining deposition and penetration levels in the downwind lower airway.
+Figure 6: Simulation-guided analytical modeling in the laryngotracheal space.
+(a) Simulated vorticity contour map on a representative two-dimensional cross-section (the S-ROAM domain) running approximately midway through the cavity; see Fig 4b.
+The core of the dominant vortex patch is conservatively bounded in the rectangle, marked by $\mathbb{V}$, with its area equating the vortex patch area $A$ used in the S-ROAM.
+Panels (b-h) show two sample particle pathlines (in red) against the S-ROAM streamlines (in grey), respectively for particle sizes 5, 10, 15, 20, 25, 30, and 50 $\mathrm{\mu m}$.
 The black lines on either side of the S-ROAM domain mark the slip wall boundaries.
-In each case, the vortex patch $\mathbb{V}$ is mimicked by a set of five point vortices embedded on a the two-dimensional flow field with background unidirectional flow in the $\beta$-direction.
-The representative particle entering near the middle of the inlet face is marked as $\mathbb{I}$ and the particle entering near the right edge of the model channel is marked as $\mathbb{II}$.
-III.4 Analytical projections for particulate transport: consistent with the full-scale numerical findings
-Figure 6a demonstrates the vorticity field mapped over the section shown in Fig.
-3c.
-In the reduced-order analytical setup, the vortex patch marked by $\mathbb{V}$ is modeled with a spatial assembly of five point vortices, as described in §II.2.
-In panels b-h of Fig.
-6, the grey curves show the streamlines of the background flow field in the S-ROAM, with the red curves tracing the sample particulate trajectories.
-Smaller particulates, owing to their low inertia and smaller $St$, follow the streamlines (on which they were embedded at entry points) more closely.
-For larger particulates, the inertia-dominated dynamics would inhibit downwind penetration through biasing their motion toward wall deposition, e.g., see panels g and (especially) h.
-Note that the particulates entering through the middle of the inlet face of the S-ROAM are marked as $\mathbb{I}$ while those entering near the right edge are marked as $\mathbb{II}$.
-As shown in Fig.
-7, the particulates $\mathbb{II}$, being nearer to the vortex region, are deviated more (from the airflow streamlines they were embedded on at the channel inlet) compared to the particulates $\mathbb{I}$.
-The deviation generically grows as the particulate diameters are increased, which aligns with the findings in Table 1.
-A greater deviation implies that the particulates are being increasingly shifted toward the channel walls, which in turn would result in a declining deposition and penetration levels in the lower airway.
-Figure 7: Comparison of the absolute deviation ($\tilde{\mathbb{D}}$, normalized with respect to the S-ROAM channel width) of the particle pathlines from the respective streamline they were embedded on at the inlet face.
-The normalized deviation is measured along the $\alpha$-direction (see Fig.
-6).
-Deviation curves for particles entering near the middle of the S-ROAM inlet face is marked as $\mathbb{I}$ and for the particles entering near the right edge of the model channel is marked as $\mathbb{II}$.
-IV Open questions: Perspectives on enhancing the biophysical realism of the modeling approach
+In each case, the vortex patch $\mathbb{V}$ is mimicked by a set of five point vortices embedded on the two-dimensional flow field with background unidirectional flow in the $\beta$-direction.
+The representative particles entering near the middle of the inlet face and near the right edge of the model channel are respectively marked as $\mathbb{I}$ and $\mathbb{II}$.
+Figure 7: Particle transport trend in the analytical case.
+Comparison of the absolute deviation ($\tilde{\mathbb{D}}$, normalized with respect to the S-ROAM channel width) of the particle pathlines from the respective streamline they were embedded on at the inlet face.
+The deviation is measured along the $\alpha$-direction (see Fig 6).
+Deviation curves for particles entering near the middle and the right edge of the model channel inlet are respectively marked as $\mathbb{I}$ and $\mathbb{II}$.
+IV Discussion: perspectives on enhancing the biophysical realism of the modeling approach
 IV.1 On the interfacial mechanics at mucociliary layers during inhalation
-While the numerical and analytical models presented here conclude that aerosols/droplets of the length scale $\mathcal{O}$($10^{1}$) $\mathrm{\mu m}$, formed through shearing and disintegration of intra-URT mucosal filaments, can have significant deposition and penetration in the deep lungs (e.g., see Table 1); to enhance the physiological realism of the model—it is essential that we obtain the actual size distribution and formation rate of liquid particulates generated through viscoelastic separation as the inhaled air brushes past the URT mucus.
-The present study employed the conservative estimate that 1 particulate of each tested size is formed during each breathing cycle and is a limitation.
-The reader should however note that unlike expulsion regimes (during exhalation), where imaging-based data collection is relatively straightforward with human subjects expelling particulates into the outside air for different speech and breathing parameters, the current problem of characterizing the internal reverse transport (inhalation) into the lower airway could be somewhat challenging.
-Figure 8: (a) Nasopharynx (compare with labels on Fig.
-1), which is the initial trigger site for SARS infection owing to the propensity of ciliated epithelial cells with angiotensin-converting enzyme 2 (ACE2) surface receptors that the spike protein viruses bind to during cell intrusion.
-(b) Representative planar outline of the surface topology at the nasopharynx.
-Phase 1 stands in for the inhaled air flux, phase 2 signifies the mucus film.
-The blue arrow is a cartooned depiction of the inhaled airflow profile sweeping over the mucosal topography (in pale red).
-The black curve in panel (b) delineating the mucosal surface topography is extracted from CT data.
-With in vivo tests in human subjects all but unlikely to receive institutional review board approvals, the approach could be essentially two-fold with synergistic numerical modeling and experimental visualizations.
-One can consider a 2-phase interaction on anatomically realistic upper airway surfaces; the two phases being the mimicked versions of inhaled air (phase 1) and the relatively static mucosal film (phase 2); see Fig.
-8.
-The interfacial mechanics leading to particulate formation and release, along with their spatiotemporal growth and size distribution, could then be assessed with multiphase simulation techniques.
-While similar problems in a different context have been studied before—the three-dimensional wavy anatomic topology of the URT surface can potentially reveal interesting nuances related to the dynamics.
-Experimental visualizations with air flux brushing past a viscoelastic film mimicking mucus rheology should then be able to validate the micro-scale mechanics revealed by the computational approach.
+While the numerical and analytical models presented here conclude that particulates of the length scale $\mathcal{O}$($10^{1}$) $\mathrm{\mu m}$, formed through shearing and disintegration of intra-URT mucosal filaments, can indeed have significant deposition and penetration along the bronchial tubes and in the deeper lung recesses (see Table 1); to enhance the physiological realism of the model—it is essential that we obtain the actual size distribution and formation rate of the liquid particulates generated through viscoelastic separation as the inhaled air brushes past the URT mucus.
+The present study used the conservative estimate that 1 particulate of each size tested is formed during each breathing cycle and is a limitation.
+The reader should however note that unlike expulsion regimes (during exhalation), where imaging-based data collection is relatively straightforward with human subjects expelling particulates into the outside air for different speech and breathing parameters, the current problem of characterizing the internal reverse transport (inhalation) into the lower airway in live subjects could be somewhat challenging.
+The approach could consequently be two-fold with synergistic numerical modeling and experimental visualizations.
+One can consider a 2-phase interaction on anatomically realistic upper airway surfaces; the two phases being the mimicked versions of inhaled air (phase 1) and the relatively static mucosal substrate (phase 2); see Fig 8a, and explore the interfacial mechanics leading to particulate formation and release, along with their generation rate, spatiotemporal growth, and size distribution.
 IV.2 On experimental validation of deep lung penetration through in vitro physical tests
-To verify the numerical modeling of the overall spatial transport, experiments could be conducted in 3D-printed anatomically realistic airway casts that are internally coated with (say) concentrated glycerol (with tuned dilution levels) standing in for mucus.
+To verify the numerical modeling of the overall spatial transport, further experiments could be conducted within 3D-printed anatomically realistic airway casts internally coated with (say) concentrated glycerol (with tuned dilution levels) standing in for mucus.
 Controlled air flux, with embedded aerosols, could be passed through the cast via suction mechanism set up with a vacuum pump.
-One approach (amongst others) for getting highly-resolved data on the lower airway penetration could be to use the sophisticated gamma scintigraphy technique [41], wherein the aerosolizing solution to be administered into the three-dimensional cast (with the incoming air) would be seeded with a mildly radioactive element (e.g., Technetium).
-After the particulates have landed along the airway walls, the radioactive signals (emitted by the deposited mass and hence presenting a measure of local penetration) could be compared with the in silico deposition patterns along the bronchial pathways.
-Note that for this approach to be functionally consistent, the printed cast should be separable into (at least) two parts along the sagittal mid-section to facilitate controlled liquid coating of its internal surfaces.
-The separable 3D-printed pieces will also logistically facilitate radio-signal imaging of the wall deposits.
-V Conclusion: the main inferences
+One approach (among others) to obtain highly resolved data on lower airway penetration could be to use the sophisticated gamma scintigraphy technique [64], wherein the solution to be aerosolized and administered into the cast (with the incoming air) would be seeded with a mildly radioactive element (e.g., Technetium).
+After the particulates have landed along the airway walls, the radioactive signals (emitted by the deposited mass and thus presenting a measure of local penetration) could be compared with the in silico deposition patterns along the bronchial pathways.
+In this context, the reader may scan our previous works with experimental validations [64, 65]; in particular, [64] called on gamma scintigraphy measurements within anatomically accurate and transparent 3D-printed airway casts to physically verify numerically modeled sinonasal deposition from over-the-counter nasal sprays.
+The modeling protocol therein implemented similar computational schemes as in the present analysis.
+IV.3 On limitations implicit to the modeling framework and its clinical relevance
+In terms of achieving true biological realism, it is important to note the limitations implicit to the numerical modeling and theoretical analysis described in this study.
+The approach here algorithmizes the complex dynamics of infection onset, particularly omitting the role of immune responses and mucosal properties.
+In an actual infection setting, the innate and adaptive immune responses would be activated, likely leading to dynamic alterations in mucus properties that could impact intra-URT particulate formation, clearance, and the pathogen’s state within these fragments.
+By focusing on a static view of mucosal fragmentation and particulate generation, this study currently does not account for such dynamic physiological responses that may influence the outcomes of viral transport and downwind deposition.
+Figure 8: Interfacial interactions and morphological complexity along mucus-coated upper airway walls.
+(a) Representative planar outline of the nasopharyngeal surface topology.
+Phase 1 = inhaled air; Phase 2 = mucus substrate.
+The blue arrow indicates the inhaled air flux sweeping over the mucus (in pale red).
+(b) Tortuosity measurement of the anatomical space using a sagittal cutting plane.
+Next, the reported results, while establishing the basic plausibility of the URT-to-LRT transmission mechanism through the use of a test case, are for a specific anatomical geometry and baseline parameters that do not (yet) explicitly account for patient-to-patient variability.
+Realistic differences in airway anatomy, mucosal characteristics, and immune response between individuals are factors that could lead to different fragmentation and transport dynamics.
+Future work should incorporate sensitivity studies that test the robustness of the modeling approach across a range of physiological conditions and individual variability.
+Moreover, the “so what” question remains.
+The findings, while precise in a computational sense and confirming the possibility of rapid lung invasion orchestrated by URT-derived pathogen-laden particulates, do not yet translate into actionable clinical recommendations.
+Future research should assess whether the insights gained could inform clinical practices or support diagnostic tools, possibly through personalized respiratory physics models that account for specific physiological traits.
+Such models could potentially identify subjects at higher risk of clinical morbidity, e.g., from brisk onset of severe lung infection.
+V Conclusion: the main takeaways
 V.1 Can large particulates, generated from the intra-URT mucus coating during inhalation, penetrate to the bronchial airspace?
-The particulate transmission trends (see Table 1) from the numerical experiments performed in this study are found to be consistent with the first-principles reduced-order analytical findings that modeled the impact of dominant intra-airway vortex instabilities in the laryngotracheal domain on local particle transport (Fig.
-6).
-Clearly, the results establish that not only the aerosols (i.e., particulates with diameters $\lesssim$ 5 $\mathrm{\mu m}$) but also droplets as large as 10 and 15 $\mathrm{\mu m}$ can exhibit remarkable efficiency at reaching the bronchial spaces and deep lungs, so long as they are sheared away from the URT surface and are still airborne as they enter the larynx.
-This stands in contrast to the general perception that only the particulates that are smaller than 5 $\mathrm{\mu m}$ can penetrate to the lower airway.
-Missing the nuance therein is the fact that the incumbent perspective is based on the mechanics of particulates inhaled from the outside air, and our understanding of intra-URT mucosal break-up during inhalation and the subsequent aerial advection profiles in the downwind tract is still nascent.
-Figure 9: (a) Sagittal cutting plane through the anatomical airway to extract the intersected curve shown in panel (b) The latter is used to evaluate the tortuosities between P-R and Q-R tracts of the airway.
-For a physics-based rationale to explain the derived deposition and penetration profiles, let us consider the tortuosity $\mathcal{T}$ of the pathways to be traversed by a particulate inhaled from outside (called hereafter $P_{o}$), compared to a particulate that is generated from the URT and is still airborne in the laryngeal airspace (hereafter called $P_{u}$).
-Figure 9 illustrates a representative vertical cutting plane used to extract the tortuosity measurements.
+The bronchial transmission trends from the numerical experiments (see Table 1 and Fig 5) are found to be consistent with the first-principles reduced-order analytical findings that modeled the impact of dominant intra-airway vortex instabilities in the laryngotracheal domain on local particle transport (Fig 6).
+Not only the aerosols (i.e., particulates with diameters $\lesssim$ 5 $\mathrm{\mu m}$) but also droplets as large as 10 and 15 $\mathrm{\mu m}$ exhibit remarkable efficiency at reaching the bronchial spaces and deep lungs, so long as they are sheared away from the URT surface and are still airborne as they enter the larynx.
+This contradicts the general perception that only particulates smaller than 5 $\mathrm{\mu m}$ can comfortably penetrate the lower airway.
+Missing the nuance therein is the fact that the incumbent perspective is based on the mechanics of particulates inhaled from the outside air, and our understanding of intra-URT mucosal breakup during inhalation and the subsequent aerial advection in the downwind tract is still nascent.
+For a physics-based rationale to explain the derived deposition and penetration profiles, let us consider the tortuosity $\mathcal{T}$ of the pathways to be traversed by a particulate inhaled from outside (called hereafter $P_{o}$), compared to a particulate that is generated within the URT and is still airborne in the laryngeal airspace (hereafter called $P_{u}$).
+Fig 8b illustrates a representative vertical cutting plane used to extract the tortuosity measurements.
 For $P_{o}$, the corresponding tortuosity is the ratio of the curved path length and the linear distance in space between the points P and R; let us represent it as $\mathcal{T}_{P_{o}}$.
 For $P_{u}$, it is similarly the ratio of the curved path length and the linear distance in space between the points Q and R; let us represent it as $\mathcal{T}_{P_{u}}$.
-The geometric measurements of the spatial distances return: $\mathcal{T}_{P_{o}}\approx 1.90$ and $\mathcal{T}_{P_{u}}\approx 1.04$.
-Consequently, if the mechanics of $P_{o}$ is inertia-dominated (true for particulates $\gtrsim$ 5 $\mathrm{\mu m}$), they exhibit less success at navigating the highly tortuous pathway (which would have otherwise led them to the lower airway) and are deposited along the anterior URT.
-On the contrary, particulates of similar sizes, if they only have to traverse the Q-R tract (see Fig.
-7), the less tortuous pathway ensures that a high percentage of them would end up depositing along (or, penetrating to) the bronchial domains.
-As an aside, also note that the tortuosity estimates obtained here match exactly with our previously published data on mammalian airway morphology with different test airway cavities [42]; see panel (d) in Fig.
-1 of the cited article.
-Thus, the findings from this study satisfactorily address the question $Q_{1}$ posed in the introduction (§I), a variant of which is the question heading this subsection.
+The geometric measurements return: $\mathcal{T}_{P_{o}}\approx 1.90$ and $\mathcal{T}_{P_{u}}\approx 1.04$.
+Consequently, if the mechanics of $P_{o}$ is inertia-dominated (true for particulates $\gtrsim$ 5 $\mathrm{\mu m}$), they exhibit less success at navigating the highly tortuous pathway and are deposited along the anterior URT; see Fig 3 in this context.
+On the contrary, particulates of similar sizes, if they only have to traverse the Q-R tract, the less tortuous pathway ensures that a high percentage of them would end up reaching the bronchial domains; see Fig 5a-b.
+As an aside, also note that the tortuosity estimates obtained here match exactly with our previously published data on mammalian airway morphology [66]; see panel (d) in figure 1 of the cited paper.
+Thus, the findings of this study satisfactorily address the question $Q_{1}$ (see introduction).
 V.2 How do the transmitted viral loads compare to the infectious dose of the test pathogen?
 Infectious dose, $I_{D}$, of a virus quantifies the minimum number of virions that can potentially launch infection in an exposed subject [67, 49] and is a fundamental virological parameter.
-Independent studies by us [8, 44] and others [9, 45] have verified that $I_{D}\approx$ 300, for SARS-CoV-2 onset.
+Independent studies by us [25, 26] and others [24, 68] have verified that $I_{D}\approx$ 300, for SARS-CoV-2.
 Thus, evidently (per Table 2), the viral load transmitted by, e.g., the 15-$\mathrm{\mu m}$ droplets would alone exceed the $I_{D}$ threshold, thereby providing a mechanics-based rationale for the fast disease progression to the lower airway.
-The brisk pace is otherwise difficult to explain based exclusively on tissue level proliferation and direct deep lung inhalation of sub 5-$\mathrm{\mu m}$ particulates from outside.
-However, from a translational paradigm, it is critical to note that the present in silico framework does not yet take into account the host innate and adaptive immune responses to the invading virions.
-The immunological considerations, once incorporated into the mechanics paradigm, can help rationalize the varying rates of clinical prognosis recorded in different subjects [38]; e.g., the deep lung infection for SARS-CoV-2 has historically ensued over a range of $\mathrm{2-8}$ days in different subjects following the appearance of initial symptoms along the upper respiratory tract.
-In summary, the findings do satisfactorily address the question $Q_{2}$ posed in §I, a variant of which is the question heading this subsection.
+The brisk pace is otherwise difficult to explain based exclusively on tissue level proliferation and direct deep lung inhalation of dominantly sub 5-$\mathrm{\mu m}$ particulates from outside.
+However, from a translational perspective, it is critical to note that the present in silico framework does not yet take into account the host innate and adaptive immune responses to the invading virions which can lead to shifts in mucus properties altering breakup, while affecting the dynamics of clearance and state of the virus in the URT-derived fragments.
+The immunological considerations, once incorporated into the mechanics paradigm, can help rationalize the varying rates of clinical prognosis recorded in different subjects [59]; e.g., deep lung infection for SARS-CoV-2 has historically ensued over a range of $\mathrm{2-8}$ days following the appearance of initial URT symptoms.
+In summary, the findings do satisfactorily address the question $Q_{2}$ posed in the introduction.
 Acknowledgments